### v2

# Descent Methods for Vector Optimization Problems: A Majorization-Minimization Perspective

Abstract
In this paper, we develop a unified framework and convergence analysis of first-order methods for vector optimization problems (VOPs) from a majorization-minimization perspective. By selecting different surrogate functions, the general method can be reduced to various existing first-order methods, both with and without line search. Our unified convergence analysis reveals that the slow convergence of the steepest descent method is primarily attributed to the significant gap between the surrogate and objective functions. Consequently, narrowing this surrogate gap can enhance the performance of first-order methods. Interestingly, we elucidate that selecting a tighter surrogate function is equivalent to using an appropriate base of the dual cone in the direction-finding subproblem. Building on this insight, we employ the Barzilai-Borwein method to narrow the surrogate gap and propose a Barzilai-Borwein descent method for VOPs (BBDVO) with polyhedral cones. By reformulating the subproblem, we provide a novel perspective on the Barzilai-Borwein descent method, bridging the gap between this method and the steepest descent method. Furthermore, by selecting an appropriate base and reformulating the subproblem in the Newton method, we also provide a novel perspective on the preconditioned Barzilai-Borwein method. Finally, several numerical experiments are presented to validate the efficiency of the BBDVO.
keywords:
Multiple objective programming , Majorization-minimization optimization , Barzilai-Borwein method , Convergence rates , Polyhedral cone
MSC:
[2010] 90C29, 90C30
journal: Journal of  Temlates
1 Introduction
This paper focuses on the following vector optimization problem:
${\rm min}_{K}~{}F(x),$ (VOP)
where $F:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ is to be optimized under the partial order induced by a closed, convex, and pointed cone $K\subset\mathbb{R}^{m}$ with a non-empty interior, defined as follows:
$y\preceq_{K}(\mathrm{resp.}\prec_{K})y^{\prime}\ \Leftrightarrow\ y^{\prime}-y\in K(\mathrm{resp.}\ \mathrm{int}(K)).$
Let $K^{*}=\{c^{*}\in\mathbb{R}^{m}:\langle c^{*},y\rangle\geq 0,\forall y\in K\}$ be the positive polar cone of $K$, and $C$ be a compact convex set such that $0\notin C$ and $\mathrm{cone}(C)=K^{*}$, namely, $C$ is a base of $K^{*}$. In vector optimization, it is often impossible to improve all objectives simultaneously with respect to the partial order. Therefore, the concept of optimality is defined as efficiency [25], meaning that there is no better solution for an efficient solution. Specifically, the problem (VOP) corresponds to a multiobjective optimization problem when $K=\mathbb{R}^{m}_{+}$, where +$\mathbb{R}^{m}_{+}$ denotes the non-negative orthant of $\mathbb{R}^{m}$. Various applications of multiobjective optimization problems (MOPs) can be found in engineering [35], economics [45, 20], management science [15], environmental analysis [32], machine learning [41, 47], etc. Although many real-world problems reformulated as vector-valued problems adhere to the partial order induced by +$\mathbb{R}^{m}_{+}$, some applications, such as portfolio selection in securities markets [1, 2], require partial orders induced by closed convex cones other than the non-negative orthant. Consequently, vector optimization problems (VOPs) have garnered significant attention in recent years.
Over the past two decades, descent methods have received increasing attention within the multiobjective optimization community, primarily due to the seminal work on the steepest descent method proposed by Fliege & Svaiter [17]. Inspired by Fliege and Svaiter’s contributions, researchers have extended other numerical algorithms to solve multiobjective optimization problems (MOPs) [see, e.g., 16, 40, 38, 18, 5, 36, 37, 43]. To the best of our knowledge, the study of descent methods for unconstrained vector optimization problems can be traced back to the work of Graña Drummond & Svaiter [14], who extended the steepest descent method for MOPs (SDMO) proposed by Fliege & Svaiter [17] to VOPs. In this context, the direction-finding subproblem at $x^{k}$ is formulated as follows:
$\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{c^{*}\in C}\left\langle{c^{*},JF(x^{k})d}\right\rangle+\frac{1}{2}\|d\|^{2},$
where $JF(x^{k})\in\mathbb{R}^{m\times n}$ is the Jacobian matrix of $F(\cdot)$ at $x^{k}$. Similar to MOPs, several standard numerical algorithms have been extended to VOPs, including the Newton method [13], projected gradient method [22], proximal point method [4], conjugated gradient method [33] and conditional gradient method [10].
In recent years, complexity analysis of descent methods for MOPs has been extensively studied. Fliege et al. [19] established the convergence rates of SDMO under different convexity assumptions. Tanabe et al. [44] developed convergence results for the multiobjective proximal gradient method. Additionally, Lapucci [31] studied the complexity of a wide class of multiobjective descent methods with nonconvex assumption. However, Chen et al. [8] noted that both theoretical and empirical results indicate that existing multiobjective first-order methods exhibit slow convergence due to objective imbalances. To address this challenge, Chen et al. [7] proposed a Barzilai-Borwein descent method for MOPs (BBDMO) that dynamically tunes gradient magnitudes using Barzilai-Borwein’s rule [3] in direction-finding subproblem. From a theoretical perspective, an improved linear convergence rate is confirmed by Chen et al. [8], demonstrating that Barzilai-Borwein descent methods can effectively mitigate objective imbalances. Despite the extensive study of the complexity analysis of descent methods for MOPs, corresponding results for VOPs have received little attention. As described by Chen et al. [6], the linear convergence rates of first-order descent methods for VOPs are influenced by $C$, the base of $K^{*}$ in direction-finding subproblem. This naturally leads to the following question:
(Q1) How to select an appropriate base to accelerate convergence ?
Similarly, for the subproblem of SDMO:
$\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{\lambda\in\Delta_{m}}\left\langle{\lambda,JF(x^{k})d}\right\rangle+\frac{1}{2}\|d\|^{2},$
where $\Delta_{m}:=\{\lambda\succeq 0:\sum_{i=1}^{m}\lambda_{i}=1\}$ is a base of +$\mathbb{R}^{m}_{+}$, a natural question arises:
(Q2) Is $\Delta_{m}$ a good choice of base for +$\mathbb{R}^{m}_{+}$ in SDMO? If not, how to choose a better one?
The majorization-minimization principle is a versatile tool for designing novel algorithms, which has been successful employed in nonconvex optimization [30], constrained optimization [29], and incremental optimization [34, 27]. The core idea behind the majorization-minimization method is to iteratively minimize a sequence of easily solvable upper bounds of the objective function. In this paper, we extend majorization-minimization method to vector optimization, addressing the aforementioned questions. Our work aims to provide a unified framework and convergence analysis of descent methods for VOPs from a majorization-minimization perspective. The primary contributions of this paper are summarized as follows:
(i) Before presenting the majorization-minimization method and its convergence analysis, we first define the concepts of strong convexity and smoothness for a vector-valued function with respect to a partial order. Leveraging these properties, we extend the notion of the condition number to VOPs, which plays a pivotal role in establishing the linear convergence of first-order methods for VOPs.
(ii) We devise a generic majorization-minimization descent method for VOPs and develop its convergence analysis. By selecting different surrogate functions, the general method can be reduced to several existing descent methods. We emphasize that the gap between the surrogate and objective functions significantly affects the performance of descent methods, which plays a central role in majorization-minimization optimization. Specifically, the steepest descent method for VOPs exhibits slow convergence due to the large gap between the surrogate and objective functions. To address this issue, we develop an improved descent method with a tighter surrogate function, resulting in improved linear convergence, and the rate of convergence is determined by the condition number. Interestingly, we show that selecting a tighter surrogate function is equivalent to using an appropriate base in the direction-find subproblem (see Remark 4.4). This provides a positive answer to (Q1). Especially, theoretical results suggest a tighter surrogate function by using Barzilai-Borwein method.
(iii) We propose a Barzilai-Borwein descent method for VOPs (BBDVO) with polyhedral cones. By reformulating the subproblem, we observe that BBDMO is essentially the SDMO with an appropriately chosen base in the direction-finding subproblem (see Remark 6.3). This observation provides a positive answer to (Q2) and offers new insights into BBDMO from a majorization-minimization perspective. Notably, a VOP with a polyhedral cone can be transformed into an MOP using a transform matrix, which is often used to define a specific polyhedral cone. We demonstrate that the performance of BBDVO is insensitive to the choice of the transform matrix, whereas the steepest descent method is highly sensitive to it. Numerical results also confirm that BBDVO is an efficient method for solving VOPs with polyhedral cones.
(iv) The fast convergence of Newton method for MOPs [16] comes at a computationally expensive per-step cost. By selecting an appropriate base and reformulating the subproblem in the Newton method, we propose an efficient second-order method for MOPs that coincides with the preconditioned Barzilai-Borwein method for MOPs (PBBMO) [9]. This formulation offers novel insights into PBBMO and highlights the indispensability of the Barzilai-Borwein rule in the quasi-Newton method from a majorization-minimization perspective.
The remainder of this paper is organized as follows. Section 2 introduces the necessary notations and definitions for later use. In Section 3, we present a generic majorization-minimization descent method for VOPs and analyze its convergence rates under various convexity assumptions. Sections 4 and 5 explore the connections between different descent methods from the perspective of majorization-minimization. Section 6 proposes a Barzilai-Borwein descent method specifically designed for VOPs with a polyhedral cone. In Section 7, we study second-order methods from a majorization-minimization perspective. Section 8 provides numerical results to demonstrate the efficiency of the BBDVO. Finally, conclusions are presented at the end of the paper.
2 Preliminaries
[…]
2.1 Vector optimization
In the subsection, we revisit some definitions and results pertinent to VOPs. Firstly, we introduce the concept of efficiency.
Definition 2.1.
[25] A vector $x^{*}\in\mathbb{R}^{n}$ is called efficient solution to (VOP) if there exists no $x\in\mathbb{R}^{n}$ such that $F(x)\preceq_{K}F(x^{\ast})$ and $F(x)\neq F(x^{\ast})$.
Definition 2.2.
[25] A vector $x^{*}\in\mathbb{R}^{n}$ is called weakly efficient solution to (VOP) if there exists no $x\in\mathbb{R}^{n}$ such that $F(x)\prec_{K}F(x^{\ast})$.
Definition 2.3.
[14] A vector $x^{\ast}\in\mathbb{R}^{n}$ is called $K$-stationary point to (VOP) if
$\mathrm{range}(JF(x^{*}))\cap(-\mathrm{int}(K))=\emptyset,$
where $\mathrm{range}(JF(x^{*}))$ denotes the range of linear mapping given by the matrix $JF(x^{*})$.
Definition 2.4.
[14] A vector $d\in\mathbb{R}^{n}$ is called $K$-descent direction for $F(\cdot)$ at $x$ if
$JF(x)d\in-\mathrm{int}(K).$
Remark 2.1.
Note that if $x\in\mathbb{R}^{n}$ is a non-stationary point, then there exists a $K$-descent direction $d\in\mathbb{R}^{n}$ such that $JF(x)d\in-\mathrm{int}(K)$.
Next, we introduce the concept of $K$-convexity for $F(\cdot)$.
Definition 2.5.
[25] The objective function $F(\cdot)$ is called $K$-convex if
$F(\lambda x+(1-\lambda)y)\preceq_{K}\lambda F(x)+(1-\lambda)F(y),\ \forall x,y\in\mathbb{R}^{n},\ \lambda\in[0,1].$
By the continuous differentiability of $F(\cdot)$, $K$-convexity of $F(\cdot)$ is equivalent to
$JF(x)(y-x)\preceq_{K}F(y)-F(x),\ \forall x,y\in\mathbb{R}^{n}.$
We conclude this section by elucidating the relationship between $K$-stationary points and weakly efficient solutions.
Lemma 2.1.
[25] Assume that the objective function $F(\cdot)$ is $K$-convex, then $x^{*}\in\mathbb{R}^{n}$ is a $K$-stationary point of (VOP) if and only if $x^{*}$ is a weakly efficient solution of (VOP).
2.2 Strong convexity and smoothness
Strong convexity and smoothness of objective functions play a central role of first-order methods in optimization. This subsection is devoted to strong convexity and smoothness of vector-valued functions under partial order.
Definition 2.6.
[13] The objective function $F(\cdot)$ is called strongly $K$-convex with $\bm{\mu}\in K$ if
$F(\lambda x+(1-\lambda)y)\preceq_{K}\lambda F(x)+(1-\lambda)F(y)-\frac{1}{2}\lambda(1-\lambda)\left\lVert{x-y}\right\rVert^{2}\bm{\mu},\ \forall x,y\in\mathbb{R}^{n},\ \lambda\in[0,1],$
and the above relation does not hold for any $\hat{\bm{\mu}}$ with $\hat{\bm{\mu}}\not\preceq_{K}\bm{\mu}$.
Remark 2.2.
Comparing with the definition in [13], Definition 2.6 includes the case $\bm{\mu}\in{\rm{bd}}(K)$, then it reduces $K$-convexity when $\bm{\mu}=0$. Furthermore, tthe final statement in the definition establishes the uniqueness of the parameter $\bm{\mu}$, which is essential for the convergence analysis.
Lemma 2.2.
[25] Assume that the objective function $F(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$, then $x^{*}\in\mathbb{R}^{n}$ is a $K$-stationary point of (VOP) if and only if $x^{*}$ is an efficient solution of (VOP).
By the continuous differentiability of $F(\cdot)$, strong $K$-convexity of $F(\cdot)$ is equivalent to
$\frac{1}{2}\left\lVert{x-y}\right\rVert^{2}\bm{\mu}+JF(x)(y-x)\preceq_{K}F(y)-F(x),\ \forall x,y\in\mathbb{R}^{n},$
it characterizes a quadratic lower-bound of $F(\cdot)$. Intuitively, we use quadratic upper-bound to define the $K$-smoothness of $F(\cdot)$ under partial order.
Definition 2.7.
The objective function $F(\cdot)$ is called $K$-smooth with $\bm{\ell}\in{\rm int}(K)$ if
$F(y)-F(x)\preceq_{K}JF(x)(y-x)+\frac{1}{2}\left\lVert{x-y}\right\rVert^{2}\bm{\ell},\ \forall x,y\in\mathbb{R}^{n},$
and the above relation does not hold for any $\hat{\bm{\ell}}$ with $\bm{\ell}\not\preceq_{K}\hat{\bm{\ell}}$.
Remark 2.3.
Assume that $F(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in K$ and $K$-smooth with $\bm{\ell}\in{\rm int}(K)$, then $\bm{\mu}\preceq_{K}\bm{\ell}$.
Remark 2.4.
Comparing with the smoothness and strong convexity in [6, Definitions 7 and 8] with Euclidean distance, i.e., $\omega(\cdot)=\frac{1}{2}\left\lVert{\cdot}\right\rVert$, Definitions 2.6 and 2.7 are tighter and do not depend on the reference vector $e$.
Next, we characterize the properties of the difference of two functions.
Lemma 2.3 (regularity of residual functions).
Let $F,G:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ be two vector-valued functions. Define $H(\cdot):=G(\cdot)-F(\cdot)$. Then the following statements hold.
  • (i)
    if $G(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$ and $F(\cdot)$ is $K$-smooth with $\bm{\ell}\in{\rm int}(K)$, where $\bm{\ell}\preceq\bm{\mu}$, then $H(\cdot)$ is strongly $K$-convex with $\bm{\mu}-\bm{\ell}$;
  • (ii)
    if $G(\cdot)$ is $K$-smooth with $\bm{\ell}\in{\rm int}(K)$ and $F(\cdot)$ is $K$-convex, then $H(\cdot)$ is $K$-smooth with $\bm{\ell}\in{\rm int}(K)$.
  • (iii)
    if $G(\cdot)$ $K$-smooth with $\bm{\ell}\in{\rm int}(K)$ and $F(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$, where $\bm{\mu}\preceq\bm{\ell}$, then $H(\cdot)$ is $K$-smooth with $\bm{\ell}-\bm{\mu}$.
Proof.
The proof is a consequence of the definition of strong $K$-convexity and $K$-smoothness, we omit it here. ∎
In SOPs, the condition number (the quotient of smoothness parameter and the modulus of strong convexity) plays a key role in the geometric convergence of first-order methods. We end this section with the definition of the condition number of a strongly $K$-convex function under partial order.
Definition 2.8.
Assume that $F(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$ and $K$-smooth with $\bm{\ell}\in{\rm int}(K)$. Then, we denote
$\kappa_{F,\preceq_{K}}:=\max_{c^{*}\in C}\frac{\left\langle{c^{*},\bm{\ell}}\right\rangle}{\left\langle{c^{*},\bm{\mu}}\right\rangle}$ (1)
the condition number of $F(\cdot)$ under partial order $\preceq_{K}$.
Remark 2.5.
Notice that $0\notin C$ and $K^{*}=\mathrm{cone}(C)$, the condition number can be rewritten as follows:
$\kappa_{F,\preceq_{K}}:=\max_{c^{*}\in K^{*}\setminus\{0\}}\frac{\left\langle{c^{*},\bm{\ell}}\right\rangle}{\left\langle{c^{*},\bm{\mu}}\right\rangle}.$
In other words, the condition number of $F(\cdot)$ is determined only by $K$, not $C$.
In the following, we will show that the condition number can be reduced to that of MOPs [8].
Proposition 2.1.
For any $\bm{\ell},\bm{\mu}\in\mathbb{R}^{m}_{++}$, we have
$\max\limits_{\lambda\in\Delta_{m}}\frac{\sum_{i\in[m]}\lambda_{i}\bm{\ell}_{i}}{\sum_{i\in[m]}\lambda_{i}\bm{\mu}_{i}}=\max\limits_{i\in[m]}\frac{\bm{\ell}_{i}}{\bm{\mu}_{i}}.$
Proof.
Without loss of generality, we assume that
$\frac{\bm{\ell}_{1}}{\bm{\mu}_{1}}\geq\frac{\bm{\ell}_{2}}{\bm{\mu}_{2}}\geq\cdots\geq\frac{\bm{\ell}_{m}}{\bm{\mu}_{m}}.$ (2)
Then, we infer that
$\max\limits_{\lambda\in\Delta_{m}}\frac{\sum_{i\in[m]}\lambda_{i}\bm{\ell}_{i}}{\sum_{i\in[m]}\lambda_{i}\bm{\mu}_{i}}\geq\frac{\bm{\ell}_{1}}{\bm{\mu}_{1}}=\max\limits_{i\in[m]}\frac{\bm{\ell}_{i}}{\bm{\mu}_{i}}.$
In the following, we need to prove
$\max\limits_{i\in[m]}\frac{\bm{\ell}_{i}}{\bm{\mu}_{i}}\geq\max\limits_{\lambda\in\Delta_{m}}\frac{\sum_{i\in[m]}\lambda_{i}\bm{\ell}_{i}}{\sum_{i\in[m]}\lambda_{i}\bm{\mu}_{i}}.$
It is sufficient to prove that
$\frac{\bm{\ell}_{1}}{\bm{\mu}_{1}}\geq\frac{\sum_{i\in[m]}\lambda_{i}\bm{\ell}_{i}}{\sum_{i\in[m]}\lambda_{i}\bm{\mu}_{i}}$
holds for all $\lambda\in\Delta_{m}.$ Let us first state the fact that for any $a_{1}/b_{1}\geq a_{2}/b_{2}>0,c,d>0$, we have
$\frac{a_{1}}{b_{1}}\geq\frac{ca_{1}+da_{2}}{cb_{1}+db_{2}}\geq\frac{a_{2}}{b_{2}}.$ (3)
For any $\lambda\in\Delta_{m}$, we denote $\mathcal{I}(\lambda):=\{i:\lambda_{i}>0\}$ with element $i_{1}<i_{2}<\cdots<i_{|\mathcal{I}(\lambda)|}$. Substituting $a_{1}/b_{1}={\bm{\ell}_{i_{1}}}/{\bm{\mu}_{i_{1}}}$, $a_{2}/b_{2}={\bm{\ell}_{i_{2}}}/{\bm{\mu}_{i_{2}}}$, $c=\lambda_{i_{1}}$ and $d=\lambda_{i_{2}}$ into (3) yields
$\frac{\bm{\ell}_{i_{1}}}{\bm{\mu}_{i_{1}}}\geq\frac{\lambda_{i_{1}}\bm{\ell}_{i_{1}}+\lambda_{i_{2}}\bm{\ell}_{i_{2}}}{\lambda_{i_{1}}\bm{\mu}_{i_{1}}+\lambda_{i_{2}}\bm{\mu}_{i_{2}}}\geq\frac{\bm{\ell}_{i_{2}}}{\bm{\mu}_{i_{2}}}.$
For $2\leq j\leq|\mathcal{I}(\lambda)|$, by substituting $a_{1}/b_{1}={\sum_{s\in[j-1]}\lambda_{i_{s}}\bm{\ell}_{i_{s}}}/{\sum_{s\in[j-1]}\lambda_{i_{s}}\bm{\mu}_{i_{s}}}$, $a_{2}/b_{2}={\bm{\ell}_{i_{j}}}/{\bm{\mu}_{i_{j}}}$, $c=1$ and $d=\lambda_{i_{j}}$ into (3) recursively, we have
$\frac{\bm{\ell}_{i_{1}}}{\bm{\mu}_{i_{1}}}\geq\cdots\geq\frac{\sum_{s\in[j-1]}\lambda_{i_{s}}\bm{\ell}_{i_{s}}}{\sum_{s\in[j-1]}\lambda_{i_{s}}\bm{\mu}_{i_{s}}}\geq\cdots\geq\frac{\sum_{s\in[|\mathcal{I}(\lambda)|]}\lambda_{i_{s}}\bm{\ell}_{i_{s}}}{\sum_{s\in[|\mathcal{I}(\lambda)|]}\lambda_{i_{s}}\bm{\mu}_{i_{s}}}.$
This, together with the definition of $\mathcal{I}(\lambda)$ and the relation (2), implies
$\frac{\bm{\ell}_{1}}{\bm{\mu}_{1}}\geq\frac{\sum_{i\in[m]}\lambda_{i}\bm{\ell}_{i}}{\sum_{i\in[m]}\lambda_{i}\bm{\mu}_{i}}.$
Since $\lambda$ is chosen arbitrary, the proof is complete. ∎
Remark 2.6.
If $K=\mathbb{R}^{m}_{+}$, then $\kappa_{F,\preceq_{K}}=\max_{\lambda\in\Delta_{m}}{\sum_{i\in[m]}\lambda_{i}\bm{\ell}_{i}}/{\sum_{i\in[m]}\lambda_{i}\bm{\mu}_{i}}$. By Proposition 2.1, it follows that $\kappa_{F,\preceq_{K}}=\max_{i\in[m]}{\bm{\ell}_{i}}/{\bm{\mu}_{i}}$. The condition number reduces to that of MOPs [8].
Proposition 2.2.
Assume that $F(\cdot)$ is strongly $K_{1}$-convex with $\bm{\mu}_{1}\in{\rm int}(K_{1})$ and $L_{1}$-smooth with $\bm{\ell}_{1}\in{\rm int}(K_{1})$. Then, for any pointed closed cone $K_{1}$ satisfied $K_{1}\subset K_{2}$, we have $F(\cdot)$ is strongly $K_{2}$-convex with $\bm{\mu}_{2}\in{\rm int}(K_{2})$ and $K_{2}$-smooth with $\bm{\ell}_{2}\in{\rm int}(K_{2})$. Futhermore, $\bm{\ell}_{2}\preceq_{K_{2}}\bm{\ell}_{1}$, $\bm{\mu}_{1}\preceq_{K_{2}}\bm{\mu}_{2}$, and $\kappa_{F,\preceq_{K_{2}}}\leq\kappa_{F,\preceq_{K_{1}}}$.
Proof.
The proof is a consequence of the definitions of strong $K$-convexity, $K$-smoothness and condition number, we omit it here. ∎
Remark 2.7.
Proposition 2.2 demonstrates that for a given vector-valued function, enlarging the order cone simplifies the corresponding vector optimization problem.
3 Majorization-minimization optimization for VOPs
3.1 Majorization-minimization descent method for VOPs
In this section, we present a generic majorization-minimization scheme for minimizing a vector function in the sense of descent.
Data: $x^{0}\in\mathbb{R}^{n}$
1 for $k=0,1,...$ do
2       Choose a strongly $K$-convex surrogate function $G_{k}(\cdot)$ of $F(\cdot)-F(x^{k})$ near $x^{k}$
3       Update $x^{k+1}:=\mathop{\arg\min}_{x\in\mathbb{R}^{n}}\max_{c^{*}\in C_{k}}\left\langle{c^{*},G_{k}(x)}\right\rangle$
4       if $x^{k+1}=x^{k}$ then
5             return $K$-stationary point $x^{k}$
6       end if
7
8 end for
Algorithm 1 Generic_Majorization-minimization_descent_method_for_VOPs
The surrogate function $G_{k}(\cdot)$ plays a central role in the generic majorization-minimization scheme. Intuitively, $G_{k}(\cdot)$ should well approximate $F(\cdot)-F(x^{k})$ near $x^{k}$ and the related subproblem should be easy to minimize. Therefore, we measure the approximation error by $H_{k}(\cdot):=G_{k}(\cdot)-F(\cdot)+F(x^{k})$. To characterize surrogates, we introduce a class of surrogate functions, which will be used to establish the convergence results of Algorithm 1.
Definition 3.1.
For $x^{k}\in\mathbb{R}^{n}$, we call $G_{k}(\cdot)$ a first-order surrogate function of $F(\cdot)-F(x^{k})$ near $x^{k}$ when
  • (i)
    $F(x^{k+1})-F(x^{k})\preceq_{K}G_{k}(x^{k+1})$, where $x^{k+1}$ is the minimizer of $\min_{x\in\mathbb{R}^{n}}\max_{c^{*}\in C_{k}}\left\langle{c^{*},G_{k}(x)}\right\rangle$, furthermore, when $F(\cdot)-F(x^{k})\preceq_{K}G_{k}(\cdot)$ for all $x\in\mathbb{R}^{n}$, we call $G_{k}(\cdot)$ a majorizing surrogate;
  • (ii)
    the approximation error $H_{k}(\cdot)$ is $K$-smooth with $\bm{\ell}\in{\rm int}(K)$, $H_{k}(x^{k})=0$, and $JH_{k}(x^{k})=0$.
We denote by $\mathcal{S}_{\bm{\ell},\bm{\mu}}(F,x^{k})$ the set of first-order strongly $K$-convex surrogate functions with $\bm{\mu}\in{\rm int}(K)$.
Next, we characterize the properties of first-order surrogate functions.
Lemma 3.1.
Let $G_{k}(\cdot)\in\mathcal{S}_{\bm{\ell},\bm{\mu}}(F,x^{k})$ and $x^{k+1}$ be the minimizer of $\min_{x\in\mathbb{R}^{n}}\max_{c^{*}\in C_{k}}\left\langle{c^{*},G_{k}(x)}\right\rangle$. Then, for all $x\in\mathbb{R}^{n}$, we have
  • (i)
    $H_{k}(x)\preceq_{K}\frac{1}{2}\left\lVert{x-x^{k}}\right\rVert^{2}\bm{\ell}$;
  • (ii)
    $\left\langle{c^{*}_{k},F(x^{k+1})}\right\rangle+\frac{1}{2}\left\lVert{x^{k+1}-x}\right\rVert^{2}\left\langle{c^{*}_{k},\bm{\mu}}\right\rangle\leq\left\langle{c^{*}_{k},F(x)}\right\rangle+\frac{1}{2}\left\lVert{x^{k}-x}\right\rVert^{2}\left\langle{c^{*}_{k},\bm{\ell}}\right\rangle$, where $c^{*}_{k}$ is a maximizer of $\max_{c^{*}\in C_{k}}\min_{x\in\mathbb{R}^{n}}\left\langle{c^{*},G_{k}(x)}\right\rangle$.
Proof.
Assertion (i) directly follows by the $K$-smoothness of $H_{k}(\cdot)$ and the facts that $H_{k}(x^{k})=0$ and $JH_{k}(x^{k})=0$. Next, we prove the assertion (ii). By Sion’s minimax theorem [42], by denoting $c^{*}_{k}$ a maximizer of $\max_{c^{*}\in C_{k}}\min_{x\in\mathbb{R}^{n}}\left\langle{c^{*},G_{k}(x)}\right\rangle$, and $x^{k+1}$ the minimizer of $\min_{x\in\mathbb{R}^{n}}\max_{c^{*}\in C_{k}}\left\langle{c^{*},G_{k}(x)}\right\rangle$, we have $JG_{k}(x^{k+1})^{T}c^{*}_{k}=0$. This, together with the strong $K$-convexity of $G_{k}(\cdot)$, implies that
$\left\langle{c^{*}_{k},G_{k}(x^{k+1})}\right\rangle+\frac{1}{2}\left\lVert{x^{k+1}-x}\right\rVert^{2}\left\langle{c^{*}_{k},\bm{\mu}}\right\rangle\leq\left\langle{c^{*}_{k},G_{k}(x)}\right\rangle,~{}\forall x\in\mathbb{R}^{n}.$
We thus use $F(x^{k+1})-F(x^{k})\preceq_{K}G_{k}(x^{k+1})$ to get
$\left\langle{c^{*}_{k},F(x^{k+1})-F(x^{k})}\right\rangle+\frac{1}{2}\left\lVert{x^{k+1}-x}\right\rVert^{2}\left\langle{c^{*}_{k},\bm{\mu}}\right\rangle$ $\leq\left\langle{c^{*}_{k},G_{k}(x^{k+1})}\right\rangle+\frac{1}{2}\left\lVert{x^{k+1}-x}\right\rVert^{2}\left\langle{c^{*}_{k},\bm{\mu}}\right\rangle$
$\leq\left\langle{c^{*}_{k},G_{k}(x)}\right\rangle$
$=\left\langle{c^{*}_{k},F(x)-F(x^{k})}\right\rangle+\left\langle{c^{*}_{k},H_{k}(x)}\right\rangle$
$\leq\left\langle{c^{*}_{k},F(x)-F(x^{k})}\right\rangle+\frac{1}{2}\left\lVert{x^{k}-x}\right\rVert^{2}\left\langle{c^{*}_{k},\bm{\ell}}\right\rangle.$
This completes the proof. ∎
3.2 Convergence analysis
In Algorithm 1, it can be observed that it terminates either with a $K$-stationary point in a finite number of iterations or generates an infinite sequence of non-stationary points. In the subsequent analysis, we will assume that the algorithm produces an infinite sequence of non-stationary points.
3.2.1 Global convergence
Firstly, we establish the global convergence result in the nonconvex cases, the following assumptions are required.
Assumption 3.1.
Assume the following statements hold for $F(\cdot)$ and $G_{k}(\cdot)$:
  • (i)
    The level set $\mathcal{L}_{F}(x^{0}):=\{x:F(x)\preceq_{K}F(x^{0})\}$ is bounded;
  • (ii)
    There exist two bases $C$ and $\tilde{C}$ such that
    $\max_{c^{*}\in C}\left\langle{c^{*},G_{k}(x^{k+1})}\right\rangle\leq\max_{c^{*}\in C_{k}}\left\langle{c^{*},G_{k}(x^{k+1})}\right\rangle\leq\max_{c^{*}\in\tilde{C}}\left\langle{c^{*},G_{k}(x^{k+1})}\right\rangle$
    holds for all $k\geq 0$. If $x^{k}\rightarrow x^{*}$, $G_{k}(\cdot)\in\mathcal{S}_{\bm{\ell},\bm{\mu}}(F,x^{k})$ and $\max_{c^{*}\in C_{k}}\left\langle{c^{*},G_{k}(x^{k+1})}\right\rangle\rightarrow 0$, then $x^{*}$ is a $K$-stationary point to (VOP).
Remark 3.1.
The Assumption 3.1(i) is a standard condition for nonconvex cases. Moreover, the Assumption 3.1(ii) holds for many descent methods in vector (multiobjective) optimization, such as $K$-steepest descent method, Barzilai-Bowein descent method [7] and Newton method.
We are now in a position to establish the global convergence of Algorithm 1.
Theorem 3.1.
Suppose that Assumption 3.1 holds, let $\{x^{k}\}$ be the sequence generated by Algorithm 1 with $G_{k}(\cdot)\in\mathcal{S}_{\bm{\ell},\bm{\mu}}(F,x^{k})$. Then, $\{x^{k}\}$ has at least one accumulation point and every accumulation point is a non-stationary point to (VOP).
Proof.
Since $G_{k}(\cdot)\in\mathcal{S}_{\bm{\ell},\bm{\mu}}(F,x^{k})$, we have
$F(x^{k+1})-F(x^{k})\preceq_{K}G_{k}(x^{k+1}),$ (4)
and
$G_{k}(x^{k+1})\preceq_{K}G_{k}(x^{k})=H_{k}(x^{k})=0.$
Then, we conclude that $\{F(x^{k})\}$ is decreasing under partial order $\preceq_{K}$. It follows by Assumption 3.1(i) and continuity of $F(\cdot)$ that $\{x^{k}\}$ is bounded and there exists $F^{*}$ such that
$F^{*}\preceq_{K}F(x^{k}),~{}\forall k\geq 0.$
The boundedness of $\{x^{k}\}$ indicates that $\{x^{k}\}$ has at least one accumulation point. Next, we prove that any accumulation point $x^{*}$ is a non-stationary point. By summing (4) from 00 to infinity, we have
$F^{*}-F(x^{0})\preceq_{K}\sum_{k=0}^{\infty}(F(x^{k+1})-F(x^{k}))\preceq_{K}\sum_{k=0}^{\infty}G_{k}(x^{k+1}).$
It follows that
$-\infty<\max_{c^{*}\in C}\left\langle{c^{*},F^{*}-F(x^{0})}\right\rangle\leq\max_{c^{*}\in C}\left\langle{c^{*},\sum_{k=0}^{\infty}G_{k}(x^{k+1})}\right\rangle\leq\max_{c^{*}\in C_{k}}\left\langle{c^{*},\sum_{k=0}^{\infty}G_{k}(x^{k+1})}\right\rangle\leq\sum_{k=0}^{\infty}\max_{c^{*}\in\tilde{C}}\left\langle{c^{*},G_{k}(x^{k+1})}\right\rangle.$
This, together with the fact that $G_{k}(x^{k+1})\preceq_{K}G_{k}(x^{k})=0$, implies $\max_{c^{*}\in C_{k}}\left\langle{c^{*},G_{k}(x^{k+1})}\right\rangle\rightarrow 0$. For the accumulation point $x^{*}$, there exists an infinite index set $\mathcal{K}$ such that $x^{k}\stackrel{{\scriptstyle\mathcal{K}}}{{\longrightarrow}}x^{*}$. Therefore, by Assumption 3.1(ii), we conclude that $x^{*}$ is a $K$-stationary point. ∎
3.2.2 Strong convergence
In the following, we establish the strong convergence result of Algorithm 3.1 in $K$-convex cases.
Theorem 3.2.
Suppose that Assumption 3.1 holds and $F(\cdot)$ is $K$-convex, let $\{x^{k}\}$ be the sequence generated by Algorithm 1 with $G_{k}(\cdot)\in\mathcal{S}_{\bm{\ell},\bm{\mu}}(F,x^{k})$ and $0\preceq_{K}\bm{\ell}\preceq_{K}\bm{\mu}$. Then, the following statements hold:
  • (i)
    $\{x^{k}\}$ converges to a weakly efficient solution $x^{*}$ of (VOP);
  • (ii)
    $u_{0}(x^{k})\leq\frac{\ell_{\max}R^{2}}{2k},~{}\forall k\geq 1$, where $\ell_{\max}:=\max_{c^{*}\in C_{k}}\left\langle{c^{*},\bm{\ell}}\right\rangle$, $R:=\{\left\lVert{x-y}\right\rVert:x,y\in\mathcal{L}_{F}(x^{0})\}$, and
    $u_{0}(x^{k}):=\max_{x\in\mathbb{R}^{n}}\min_{c^{*}\in C_{k}}\left\langle{c^{*},F(x^{k})-F(x)}\right\rangle$
    is a merit function in the sense of weak efficiency.
Proof.
(i) By the similar arguments in the proof of Theorem 3.1, we conclude that $\{x^{k}\}$ is bounded, and there exists a $K$-stationary point $x^{*}$ such that $F(x^{*})\preceq_{K}F(x^{k})$. Besides, the $K$-convexity of $F(\cdot)$ indicates that $x^{*}$ is a weakly efficient point. From Lemma 3.1(ii), for any $x\in\mathbb{R}^{n}$ we have
$\left\langle{c^{*}_{k},F(x^{k+1})-F(x)}\right\rangle\leq\frac{1}{2}\left\lVert{x^{k}-x}\right\rVert^{2}\left\langle{c^{*}_{k},\bm{\ell}}\right\rangle-\frac{1}{2}\left\lVert{x^{k+1}-x}\right\rVert^{2}\left\langle{c^{*}_{k},\bm{\mu}}\right\rangle.$ (5)
Substituting $x=x^{*}$ into the above inequality, we obtain
$\left\langle{c^{*}_{k},F(x^{k+1})-F(x^{*})}\right\rangle\leq\frac{1}{2}\left\lVert{x^{k}-x^{*}}\right\rVert^{2}\left\langle{c^{*}_{k},\bm{\ell}}\right\rangle-\frac{1}{2}\left\lVert{x^{k+1}-x^{*}}\right\rVert^{2}\left\langle{c^{*}_{k},\bm{\mu}}\right\rangle.$
Recall that $F(x^{*})\preceq_{K}F(x^{k})$, it follows that
$\left\lVert{x^{k+1}-x^{*}}\right\rVert^{2}\left\langle{c^{*}_{k},\bm{\mu}}\right\rangle\leq\left\lVert{x^{k}-x^{*}}\right\rVert^{2}\left\langle{c^{*}_{k},\bm{\ell}}\right\rangle.$
Furthermore, we use the fact that $0\preceq_{K}\bm{\ell}\preceq_{K}\bm{\mu}$ to get
$\left\lVert{x^{k+1}-x^{*}}\right\rVert^{2}\leq\left\lVert{x^{k}-x^{*}}\right\rVert^{2}.$
Therefore, the sequence $\{\left\lVert{x^{k}-x^{*}}\right\rVert\}$ converges. This, together with the fact that $x^{*}$ is an accumulation point of $\{x^{k}\}$, implies that $\{x^{k}\}$ converges to $x^{*}$
(ii) Since $0\preceq_{K}\bm{\ell}\preceq_{K}\bm{\mu}$, we use inequality (5) and the definition of $\ell_{\max}$ to obtain
$\left\langle{c^{*}_{k},F(x^{k+1})-F(x)}\right\rangle\leq\frac{1}{2}\left\lVert{x^{k}-x}\right\rVert^{2}\left\langle{c^{*}_{k},\bm{\ell}}\right\rangle-\frac{1}{2}\left\lVert{x^{k+1}-x}\right\rVert^{2}\left\langle{c^{*}_{k},\bm{\mu}}\right\rangle\leq\frac{\ell_{\max}}{2}\left(\left\lVert{x^{k}-x}\right\rVert^{2}-\left\lVert{x^{k+1}-x}\right\rVert^{2}\right).$
Taking the sum of the preceding inequality over 00 to $k-1$, we have
$\sum\limits_{s=0}^{k-1}\left\langle{c^{*}_{s},F(x^{s+1})-F(x)}\right\rangle\leq\frac{\ell_{\max}}{2}\left(\left\lVert{x^{0}-x}\right\rVert^{2}-\left\lVert{x^{k}-x}\right\rVert^{2}\right)\leq\frac{\ell_{\max}}{2}\left\lVert{x^{0}-x}\right\rVert^{2}.$
Notice that $F(x^{k})\preceq_{K}F(x^{s+1})$ for all $s\leq k-1$, it leads to
$\sum\limits_{s=0}^{k-1}\left\langle{c^{*}_{s},F(x^{k})-F(x)}\right\rangle\leq\frac{\ell_{\max}}{2}\left\lVert{x^{0}-x}\right\rVert^{2}.$
Denote $\hat{c}^{*}_{k}:={\sum_{s=0}^{k-1}c^{*}_{s}}/{k}$, it follows by the convexity of $C_{k}$ that $\hat{c}^{*}_{k}\in C_{k}$. Therefore, we conclude that
$\left\langle{\hat{c}^{*}_{k},F(x^{k})-F(x)}\right\rangle\leq\frac{\ell_{\max}\left\lVert{x^{0}-x}\right\rVert^{2}}{2k}.$
Select $y^{k}\in\mathop{\arg\max}_{x\in\mathbb{R}^{n}}\min_{c^{*}\in C_{k}}\left\langle{c^{*},F(x^{k})-F(x)}\right\rangle$, it holds that
$u_{0}(x^{k})$ $=\max_{x\in\mathbb{R}^{n}}\min_{c^{*}\in C_{k}}\left\langle{c^{*},F(x^{k})-F(x)}\right\rangle=\min_{c^{*}\in C_{k}}\left\langle{c^{*},F(x^{k})-F(y^{k})}\right\rangle$
$\leq\left\langle{\hat{c}^{*}_{k},F(x^{k})-F(y^{k})}\right\rangle\leq\frac{\ell_{\max}\left\lVert{x^{0}-y^{k}}\right\rVert^{2}}{2k}.$
By the definition of $y^{k}$, we deduce that $y^{k}\in\{x:F(x)\preceq_{K}F(x^{k})\}\subset\mathcal{L}_{F}(x^{0})$, which implies $\left\lVert{x^{0}-z^{k}}\right\rVert\leq R$, the desired result follows. ∎
3.2.3 Linear convergence
By further assuming that $F(\cdot)$ is strongly $K$-convex, the linear convergence result of Algorithm 1 can be derived as follows.
Theorem 3.3.
Suppose that Assumption 3.1(ii) holds $F(\cdot)$ is strongly $K$-convex, let $\{x^{k}\}$ be the sequence generated by Algorithm 1 with $G_{k}(\cdot)\in\mathcal{S}_{\bm{\ell},\bm{\mu}}(F,x^{k})$ and $0\preceq_{K}\bm{\ell}\prec_{K}\bm{\mu}$. Then, the following statements hold:
  • (i)
    $\{x^{k}\}$ converges to an efficient solution $x^{*}$ of (VOP);
  • (ii)
    $\left\lVert{x^{k+1}-x^{*}}\right\rVert\leq\sqrt{\max\limits_{c^{*}\in C_{k}}\frac{\left\langle{c^{*},\bm{\ell}}\right\rangle}{\left\langle{c^{*},\bm{\mu}}\right\rangle}}\left\lVert{x^{k}-x^{*}}\right\rVert,~{}\forall k\geq 0$.
Proof.
(i) Since $F(\cdot)$ is strongly $K$-convex, then Assumption 3.1(i) holds and every weakly efficient solution is actually an efficient solution. Therefore, assertion (i) is a consequence of Theorem 3.2(i).
(ii) By substituting $x=x^{*}$ into inequality (5), we have
$\left\langle{c^{*}_{k},F(x^{k+1})-F(x^{*})}\right\rangle\leq\frac{1}{2}\left\lVert{x^{k}-x^{*}}\right\rVert^{2}\left\langle{c^{*}_{k},\bm{\ell}}\right\rangle-\frac{1}{2}\left\lVert{x^{k+1}-x^{*}}\right\rVert^{2}\left\langle{c^{*}_{k},\bm{\mu}}\right\rangle.$
It follows by $F(x^{*})\preceq_{K}F(x^{k+1})$ that
$\left\lVert{x^{k+1}-x^{*}}\right\rVert\leq\sqrt{\frac{\left\langle{c^{*}_{k},\bm{\ell}}\right\rangle}{\left\langle{c^{*}_{k},\bm{\mu}}\right\rangle}}\left\lVert{x^{k}-x^{*}}\right\rVert.$
The desired result follows. ∎
Remark 3.2.
It seems that the convexity of $F(\cdot)$ plays no role in the proof of Theorems 3.2 and 3.3. However, it can indeed be shown that $F(\cdot)$ is necessarily $K$-convex if $\bm{\ell}=\bm{\mu}$ and strongly $K$-convex with $\bm{\mu}-\bm{\ell}$ if $\bm{\ell}\prec_{K}\bm{\mu}$. In the next section, we give some examples where such a condition holds.
Remark 3.3.
The linear convergence rate is related to $G_{k}(\cdot)$, which confirms that the rate of convergence can be improved by choosing a tighter surrogate.
4 Descent methods for VOPs without line search
In this section, we will revisit some well-known descent methods for VOPs and link them to the Algorithm 1. The main aim is to provide a unified theoretical analysis for all of them. We also develop some new algorithms with faster linear convergence.
4.1 $K$-steepest descent method for VOPs without line search
For $x\in\mathbb{R}^{n}$, recall that $d^{k}$, the $K$-steepest descent direction [14] at $x^{k}$, is defined as the optimal solution of
$\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{c^{*}\in C_{k}}\left\langle{c^{*},JF(x^{k})d}\right\rangle+\frac{1}{2}\|d\|^{2}.$ (6)
Select a vector $e\in{\rm int}(K)$, and denote $C_{e}=\{c^{*}\in K^{*}:\left\langle{c^{*},e}\right\rangle=1\}$. If we set $C_{k}=C_{e}$ in (6), then the $K$-steepest descent direction can be reformulated as the optimal solution of
$\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{c^{*}\in C_{e}}\left\langle{c^{*},JF(x^{k})d+\frac{1}{2}\|d\|^{2}e}\right\rangle.$ (7)
Remark 4.1.
If $K=\mathbb{R}^{m}_{+}$, and $C_{e}=\Delta_{m}$, then $e=(1,...,1)^{T}$ and the subproblem (7) reduces to that of steepest descent method for MOPs [17].
From now on, we assume that $F(\cdot)$ is $K$-smooth with $\bm{\ell}\in{\rm int}(K)$, denote
$L_{\max}:=\max_{c^{*}\in C_{e}}\left\langle{c^{*},\bm{\ell}}\right\rangle.$
Let us revisit the $K$-steepest descent method without line search:
Data: $x^{0}\in\mathbb{R}^{n},L\geq L_{\max}$
1 for $k=0,1,...$ do
2       Update $x^{k+1}:=\mathop{\arg\min}_{x\in\mathbb{R}^{n}}\max_{c^{*}\in C_{e}}\left\langle{c^{*},JF(x^{k})(x-x^{k})}\right\rangle+\frac{L}{2}\|x-x^{k}\|^{2}$
3       if $x^{k+1}=x^{k}$ then
4             return $K$-stationary point $x^{k}$
5       end if
6
7 end for
Algorithm 2 $K$-steepest_descent_method_without_line_search_for_VOPs
We consider the following surrogate:
$G_{k,Le}(x):=JF(x^{k})(x-x^{k})+\frac{L}{2}\left\lVert{x-x^{k}}\right\rVert^{2}e.$ (8)
Proposition 4.1.
Let $G_{k,Le}(\cdot)$ be defined as (8). Then, the following statements hold.
  • (i)
    For any $L\geq L_{\max}$, $G_{k,Le}(\cdot)$ is a majorizing surrogate of $F(\cdot)-F(x^{k})$, i.e., $F(\cdot)-F(x^{k})\preceq_{K}G_{k,Le}(\cdot)$.
  • (ii)
    If $F(\cdot)$ is $K$-convex, then $G_{k,Le}(\cdot)\in\mathcal{S}_{Le,Le}(F,x^{k})$ for all $L\geq L_{\max}$.
  • (iii)
    If $F(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$, then $G_{k,Le}(\cdot)\in\mathcal{S}_{Le-\bm{\mu},Le}(F,x^{k})$ for all $L\geq L_{\max}$.
Proof.
By the definition of $L_{\max}$, we have $\bm{\ell}\preceq_{K}L_{\max}e$, it follows from the $K$-smoothness of $F(\cdot)$ that assertion (i) holds. Notice that $G_{k,Le}(\cdot)$ is strongly $K$-convex and $K$-smooth with $Le$, and $\bm{\mu}\preceq Le$, then we obtain assertion (ii) and (iii) by Lemma 2.3 (ii) and (iii), respectively. ∎
Lemma 4.1.
Assume that $F(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$, let $\{x^{k}\}$ be the sequence generated by Algorithm 2. Then, the following statements hold:
  • (i)
    $\{x^{k}\}$ converges to an efficient solution $x^{*}$ of (VOP);
  • (ii)
    $\left\lVert{x^{k+1}-x^{*}}\right\rVert\leq\sqrt{1-{\mu_{\min}}/{L}}\left\lVert{x^{k}-x^{*}}\right\rVert,~{}\forall k\geq 0$, where $\mu_{\min}:=\min_{c^{*}\in C_{e}}\left\langle{c^{*},\bm{\mu}}\right\rangle$.
Proof.
Since $F(\cdot)$ is strongly $K$-convex, it follows that $G_{k,Le}(\cdot)\in\mathcal{S}_{Le-\bm{\mu},Le}(F,x^{k})$ and Assumption 3.1 holds in this case. By setting $C=C_{e}$, Theorem 3.3 (i) and (ii) reduce to the assertions (i) and (ii), respectively. ∎
Remark 4.2.
If $K=\mathbb{R}^{m}_{+}$ and $e=(1,...,1)^{T}$, then $C_{e}$ is a base of +$\mathbb{R}^{m}_{+}$, the convergence rate in Lemma 4.1 reduces to that of [44, Theorem 5.3] with $g(\cdot)=0$. It is worth noting that the rate of convergence is related to $e$, to the best of our knowledge, apart from $e=(1,...,1)^{T}$, it remains an open problem for the better choice of the parameter in MOPs.
4.2 Improved $K$-steepest descent method for VOPs without line search
As detailed in [8], the linear convergence result in Lemma 4.1 can be improved, this is mainly due to the large gap between $F(\cdot)-F(x^{k})$ and $G_{k,Le}(\cdot)$ from a majorization-minimization perspective. Notice that $\bm{\ell}\preceq_{K}L_{\max}e$, we denote the following tighter surrogate:
$G_{k,\bm{\ell}}(x):=JF(x^{k})(x-x^{k})+\frac{1}{2}\left\lVert{x-x^{k}}\right\rVert^{2}\bm{\ell}.$ (9)
Proposition 4.2.
Let $G_{k,\bm{\ell}}(\cdot)$ be defined as (9). Then, the following statements hold.
  • (i)
    $G_{k,\bm{\ell}}(\cdot)$ is a tight majorizing surrogate of $F(\cdot)-F(x^{k})$, i.e., $F(\cdot)-F(x^{k})\preceq_{K}G_{k,\bm{\ell}}(\cdot)$, and the relation does not hold for any $G_{k,\hat{\bm{\ell}}}(\cdot)$ such that $\bm{\ell}\not\preceq_{K}\hat{\bm{\ell}}$.
  • (ii)
    If $F(\cdot)$ is $K$-convex, then $G_{k,\bm{\ell}}(\cdot)\in\mathcal{S}_{\bm{\ell},\bm{\ell}}(F,x^{k})$.
  • (iii)
    If $F(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$, then $G_{k,\bm{\ell}}(\cdot)\in\mathcal{S}_{\bm{\ell}-\bm{\mu},\bm{\ell}}(F,x^{k})$.
Proof.
The assertions can be obtained by using the similar arguments as in the proof of Proposition 4.1. ∎
By using the tighter surrogate, we devise the following improved $K$-steepest descent method without line search.
Data: $x^{0}\in\mathbb{R}^{n}$
1 for $k=0,1,...$ do
2       Update $x^{k+1}:=\mathop{\arg\min}_{x\in\mathbb{R}^{n}}\max_{c^{*}\in C_{e}}\left\langle{c^{*},JF(x^{k})(x-x^{k})+\frac{1}{2}\|x-x^{k}\|^{2}\bm{\ell}}\right\rangle$
3       if $x^{k+1}=x^{k}$ then
4             return $K$-stationary point $x^{k}$
5       end if
6
7 end for
Algorithm 3 #1_improved_$K$-steepest_descent_method_without_line_search_for_VOPs
Lemma 4.2.
Assume that $F(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$, let $\{x^{k}\}$ be the sequence generated by Algorithm 3. Then, the following statements hold:
  • (i)
    $\{x^{k}\}$ converges to an efficient solution $x^{*}$ of (VOP);
  • (ii)
    $\left\lVert{x^{k+1}-x^{*}}\right\rVert\leq\sqrt{1-{1}/{\kappa_{F,\preceq_{K}}}}\left\lVert{x^{k}-x^{*}}\right\rVert,~{}\forall k\geq 0$.
Proof.
The assertions can be obtained by using the similar arguments as in the proof of Lemma 4.1. ∎
Remark 4.3.
If $K=\mathbb{R}^{m}_{+}$, and $e=(1,...,1)^{T}$, the convergence rate in Lemma 4.2 reduces to that of [8, Corollary 4.3] with $g(\cdot)=0$. Furthermore, $1/\kappa_{F,\preceq_{K}}\geq\mu_{\min}/L$, and the improved linear convergence does not depend on the choice of $e$.
4.3 Equivalence between Algorithms 2 and 3
As we described in Remark 4.2, the linear convergence rate of Algorithms 2 depends on $e$. In response to this issue, we attempt to give a better choice of $e$ in Algorithms 2 from a complexity perspective. Recall that the linear convergence rate of algorithm 3 does not depend on the choice of $e$. We denote
$C_{\bm{\ell}}:=\{c^{*}\in K^{*}:\left\langle{c^{*},\bm{\ell}}\right\rangle=1\}.$
In the following, we propose another improved $K$-steepest descent method for VOPs.
Data: $x^{0}\in\mathbb{R}^{n}$
1 for $k=0,1,...$ do
2       Update $x^{k+1}:=\mathop{\arg\min}_{x\in\mathbb{R}^{n}}\max_{c^{*}\in C_{\bm{\ell}}}\left\langle{c^{*},JF(x^{k})(x-x^{k})}\right\rangle+\frac{1}{2}\|x-x^{k}\|^{2}$
3       if $x^{k+1}=x^{k}$ then
4             return $K$-stationary point $x^{k}$
5       end if
6
7 end for
Algorithm 4 #2_improved_$K$-steepest_descent_method_without_line_search_for_VOPs
Notice that the subproblem in Algorithm 4 can be rewritten equivalently as:
$\min_{x\in\mathbb{R}^{n}}\max_{c^{*}\in C_{\bm{\ell}}}\left\langle{c^{*},JF(x^{k})(x-x^{k})+\frac{1}{2}\|x-x^{k}\|^{2}\bm{\ell}}\right\rangle.$
Similar to Lemma 4.2, Algorithm 4 enjoys the following improved linear convergence.
Lemma 4.3.
Assume that $F(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$, let $\{x^{k}\}$ be the sequence generated by Algorithm 4. Then, the following statements hold:
  • (i)
    $\{x^{k}\}$ converges to an efficient solution $x^{*}$ of (VOP);
  • (ii)
    $\left\lVert{x^{k+1}-x^{*}}\right\rVert\leq\sqrt{1-{1}/{\kappa_{F,\preceq_{K}}}}\left\lVert{x^{k}-x^{*}}\right\rVert,~{}\forall k\geq 0$.
Remark 4.4.
If $K=\mathbb{R}^{m}_{+}$, and $\Delta_{m}^{\bm{\ell}}:=\{c^{*}\in\mathbb{R}^{m}_{++}:\left\langle{c^{*},\bm{\ell}}\right\rangle=1\}$, the Algorithm 4 reduces to [8, Algorithm 5] with $g(\cdot)=0$. Interestingly, the relations between Algorithms 2, 3 and 4 only depends on the choice of $e$. If $e=\bm{\ell}$, Algorithms 2, 3 and 4 are equivalent. As a result, for the open problem of the better choice of $e$ in Remark 4.2, we give a partially positive answer with $e=\bm{\ell}$.
Remark 4.5.
Although Algorithms 3 and 4 both exhibit similar improved linear convergence, the computational cost of solving subproblems is generally lower in Algorithm 4. Details on solving these subproblems will be provided in Section 6.
5 Descent methods for VOPs with line search
In the previous section, the majorization-minimization optimization methods were developed using majorizing surrogate functions; however, these surrogates may be overly conservative due to reliance on global upper bounds. From the perspective of majorization-minimization, selecting a non-majorizing surrogate function could potentially enhance performance.
5.1 $K$-steepest descent method with line search
In this section, we revisit $K$-steepest descent method with line search for VOPs.
Data: $x^{0}\in\mathbb{R}^{n},\gamma\in(0,1)$
1 for $k=0,1,...$ do
2       Update $d^{k}:=\mathop{\arg\min}_{d\in\mathbb{R}^{n}}\max_{c^{*}\in C_{e}}\left\langle{c^{*},JF(x^{k})d}\right\rangle+\frac{1}{2}\|d\|^{2}$
3       if $d^{k}=0$ then
4             return $C$-stationary point $x^{k}$
5      else
6            Compute the stepsize $t_{k}\in(0,1]$ in the following way:
$t_{k}:=\max\left\{\gamma^{j}:j\in\mathbb{N},F(x^{k}+\gamma^{j}d^{k})-F(x^{k})\preceq_{K}\gamma^{j}\left(JF(x^{k})d^{k}+\frac{1}{2}\left\lVert{d^{k}}\right\rVert^{2}e\right)\right\}$
7             $x^{k+1}:=x^{k}+t_{k}d^{k}$
8       end if
9
10 end for
Algorithm 5 K-steepest_descent_method_for_VOPs
The stepsize has the following lower bound.
Proposition 5.1.
The stepsize generated in Algorithm 5 satisfies $t_{k}\geq t_{\min}:=\min\left\{\frac{\gamma}{L_{\max}},1\right\}$.
Proof.
By the line search condition in Algorithm 5, we have
$F(x^{k}+\frac{t_{k}}{\gamma}d^{k})-F(x^{k})\not\preceq_{K}\frac{t_{k}}{\gamma}\left(JF(x^{k})d^{k}+\frac{1}{2}\left\lVert{d^{k}}\right\rVert^{2}e\right).$
Then there exists $c^{*}_{1}\in C_{e}$ such that
$\left\langle{c^{*}_{1},F(x^{k}+\frac{t_{k}}{\gamma}d^{k})-F(x^{k})}\right\rangle>\left\langle{c^{*}_{1},\frac{t_{k}}{\gamma}\left(JF(x^{k})d^{k}+\frac{1}{2}\left\lVert{d^{k}}\right\rVert^{2}e\right)}\right\rangle.$ (10)
On the other hand, the $K$-smoothness of $F(\cdot)$ implies
$F(x^{k}+\frac{t_{k}}{\gamma}d^{k})-F(x^{k})\preceq_{K}\frac{t_{k}}{\gamma}JF(x^{k})d^{k}+\frac{1}{2}\left\lVert{\frac{t_{k}}{\gamma}d^{k}}\right\rVert^{2}\bm{\ell}.$
Therefore, we have
$\left\langle{c^{*}_{1},F(x^{k}+\frac{t_{k}}{\gamma}d^{k})-F(x^{k})}\right\rangle\leq\left\langle{c^{*}_{1},\frac{t_{k}}{\gamma}JF(x^{k})d^{k}}\right\rangle+\frac{1}{2}\left\lVert{\frac{t_{k}}{\gamma}d^{k}}\right\rVert^{2}\left\langle{c^{*}_{1},\bm{\ell}}\right\rangle.$
This, together with inequality (10), yields
$t_{k}\geq\frac{\gamma}{\left\langle{c^{*}_{1},\bm{\ell}}\right\rangle}.$
Then the desired result follows. ∎
We consider the following surrogate:
$G_{k,e/t_{\min}}(x):=JF(x^{k})(x-x^{k})+\frac{1}{2t_{\min}}\left\lVert{x-x^{k}}\right\rVert^{2}e.$ (11)
We can show that $G_{k,e/t_{\min}}(\cdot)$ is a non-majorizing surrogate function of $F(\cdot)-F(x^{k})$ near $x^{k}$.
Proposition 5.2.
Let $G_{k,e/t_{\min}}(\cdot)$ be defined as (11). Then, the following statements hold.
  • (i)
    $F(x^{k+1})-F(x^{k})\preceq_{K}G_{k,e/t_{\min}}(x^{k+1})$.
  • (ii)
    If $F(\cdot)$ is $K$-convex, then $G_{k,e/t_{\min}}(\cdot)\in\mathcal{S}_{e/t_{\min},e/t_{\min}}(F,x^{k})$.
  • (iii)
    If $F(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$, then $G_{k,e/t_{\min}}(\cdot)\in\mathcal{S}_{e/t_{\min}-\bm{\mu},e/t_{\min}}(F,x^{k})$.
Proof.
The assertions can be obtained by using the similar arguments as in the proof of Proposition 4.1. ∎
Lemma 5.1.
Assume that $F(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$, let $\{x^{k}\}$ be the sequence generated by Algorithm 5. Then, the following statements hold:
  • (i)
    $\{x^{k}\}$ converges to an efficient solution $x^{*}$ of (VOP);
  • (ii)
    $\left\lVert{x^{k+1}-x^{*}}\right\rVert\leq\sqrt{1-t_{\min}{\mu_{\min}}}\left\lVert{x^{k}-x^{*}}\right\rVert,~{}\forall k\geq 0$, where $\mu_{\min}:=\min_{c^{*}\in C_{e}}\left\langle{c^{*},\bm{\mu}}\right\rangle$.
Proof.
The assertions can be obtained by using the similar arguments as in the proof of Lemma 4.1. ∎
Remark 5.1.
If $K=\mathbb{R}^{m}_{+}$, and $e=(1,...,1)^{T}$, the convergence rate in Lemma 5.1(ii) reduces to that of [19, Theorem 4.2].
5.2 Generic first-order method for VOPs
Select $e_{k}\in{\rm int}(K)$, we devise the following generic first-order method:
Data: $x^{0}\in\mathbb{R}^{n},\gamma\in(0,1)$
1 for $k=0,1,...$ do
2      Select $e_{k}\in{\rm int}(K)$
3       Update $d^{k}:=\mathop{\arg\min}_{d\in\mathbb{R}^{n}}\max_{c^{*}\in C_{e}}\left\langle{c^{*},JF(x^{k})d+\frac{1}{2}\|d\|^{2}e_{k}}\right\rangle$
4       if $d^{k}=0$ then
5             return $C$-stationary point $x^{k}$
6      else
7            Compute the stepsize $t_{k}\in(0,1]$ in the following way:
$t_{k}:=\max\left\{\gamma^{j}:j\in\mathbb{N},F(x^{k}+\gamma^{j}d^{k})-F(x^{k})\preceq_{K}\gamma^{j}\left(JF(x^{k})d^{k}+\frac{1}{2}\left\lVert{d^{k}}\right\rVert^{2}e_{k}\right)\right\}$
8             $x^{k+1}:=x^{k}+t_{k}d^{k}$
9       end if
10
11 end for
Algorithm 6 Generic_first_order_method_for_VOPs
It is worth noting that we don’t specify how to select $e_{k}$ in Algorithm 6. This naturally raises the question: what role does $e_{k}$ play in determining the convergence rate?
Firstly, we derive the lower bound of stepsize in each iteration.
Proposition 5.3.
The stepsize generated in Algorithm 6 satisfies $t_{k}\geq t_{k}^{\min}:=\min\left\{\min\limits_{c^{*}\in C_{e}}\frac{\gamma\left\langle{c^{*},e_{k}}\right\rangle}{\left\langle{c^{*},\bm{\ell}}\right\rangle},1\right\}$.
Proof.
The result can be obtained by using the similar arguments as in the proof of Proposition 5.1. ∎
We consider the following surrogate:
$G_{k,e_{k}/t_{k}}(x):=JF(x^{k})(x-x^{k})+\frac{1}{2t^{\min}_{k}}\left\lVert{x-x^{k}}\right\rVert^{2}e_{k}.$ (12)
We can show that $G_{k,e_{k}/t^{\min}_{k}}(\cdot)$ is a non-majorizing surrogate function of $F(\cdot)-F(x^{k})$ near $x^{k}$.
Proposition 5.4.
Let $G_{k,e_{k}/t^{\min}_{k}}(\cdot)$ be defined as (12). Then, the following statements hold.
  • (i)
    $F(x^{k+1})-F(x^{k})\preceq_{K}G_{k,e_{k}/t^{\min}_{k}}(x^{k+1})$..
  • (ii)
    If $F(\cdot)$ is $K$-convex, then $G_{k,e_{k}/t^{\min}_{k}}(\cdot)\in\mathcal{S}_{e_{k}/t^{\min}_{k},e_{k}/t^{\min}_{k}}(F,x^{k})$.
  • (iii)
    If $F(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$, then $G_{k,e_{k}/t^{\min}_{k}}(\cdot)\in\mathcal{S}_{e_{k}/t^{\min}_{k}-\bm{\mu},e_{k}/t^{\min}_{k}}(F,x^{k})$.
Proof.
The assertions can be obtained by using the similar arguments as in the proof of Proposition 4.1. ∎
The following results show that $e_{k}$ plays a significant role in the convergence rate of Algorithm 6.
Lemma 5.2.
Assume that $F(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$, let $\{x^{k}\}$ be the sequence generated by Algorithm 6. Then, the following statements hold:
  • (i)
    $\{x^{k}\}$ converges to an efficient solution $x^{*}$ of (VOP);
  • (ii)
    $\left\lVert{x^{k+1}-x^{*}}\right\rVert\leq\sqrt{1-\min\limits_{c^{*}\in C_{e}}\frac{\left\langle{c^{*},\bm{\mu}}\right\rangle}{\left\langle{c^{*},e_{k}/t^{\min}_{k}}\right\rangle}}\left\lVert{x^{k}-x^{*}}\right\rVert,~{}\forall k\geq 0$.
  • (iii)
    If $e_{k}=e$, we have
    $\left\lVert{x^{k+1}-x^{*}}\right\rVert\leq\sqrt{1-t_{\min}{\mu_{\min}}}\left\lVert{x^{k}-x^{*}}\right\rVert,~{}\forall k\geq 0,$
    where $\mu_{\min}:=\min_{c^{*}\in C_{e}}\left\langle{c^{*},\bm{\mu}}\right\rangle$.
  • (iv)
    For any $e_{k}\in\mathrm{int}K$, we have
    $\min\limits_{c^{*}\in C_{e}}\frac{\left\langle{c^{*},\bm{\mu}}\right\rangle}{\left\langle{c^{*},e_{k}/t^{\min}_{k}}\right\rangle}\leq\frac{\gamma}{\kappa_{F,\preceq_{K}}}.$ (13)
    Moreover, the equality holds with $e_{k}=\bm{\mu}$ or $e_{k}=\bm{\ell}$. In these cases, we have
    $\left\lVert{x^{k+1}-x^{*}}\right\rVert\leq\sqrt{1-{\gamma}/{\kappa_{F,\preceq_{K}}}}\left\lVert{x^{k}-x^{*}}\right\rVert,~{}\forall k\geq 0.$ (14)
Proof.
The assertions (i) and (ii) can be obtained by using the similar arguments as in the proof of Lemma 4.1. By substituting $e_{k}=e$ into (ii), we can obtain the assertion (iii). Next, we prove the assertion (iv). Since $C_{e}$ is a compact set, there exists a vector $c_{0}^{*}\in C_{e}$ such that $1/\kappa_{F,\preceq_{K}}={\left\langle{c_{0}^{*},\bm{\mu}}\right\rangle}/{\left\langle{c_{0}^{*},\bm{\ell}}\right\rangle}.$ On the other hand, by the definition of $t_{k}^{\min}$ we can deduce
$\min\limits_{c^{*}\in C_{e}}\frac{\left\langle{c^{*},\bm{\mu}}\right\rangle}{\left\langle{c^{*},e_{k}/t^{\min}_{k}}\right\rangle}\leq\frac{\left\langle{c_{0}^{*},\bm{\mu}}\right\rangle}{\left\langle{c_{0}^{*},e_{k}}\right\rangle}\frac{\gamma\left\langle{c_{0}^{*},e_{k}}\right\rangle}{\left\langle{c_{0}^{*},\bm{\ell}}\right\rangle}=\frac{\gamma}{\kappa_{F,\preceq_{K}}}.$
Then the relation (13) follows. The equality can be obtain by substituting $e_{k}=\bm{\mu}$ or $e_{k}=\bm{\ell}$ into the left-hand side of (13) . Moreover, The equality leads to the relation (14). ∎
Remark 5.2.
As described in Lemma 5.2(iv), by setting $e_{k}=\bm{\mu}$ or $e_{k}=\bm{\ell}$, we can derive the optimal linear convergence rate for Algorithm 6. Moreover, the convergence rate reduces to that of Lemma 4.3(ii) with constant $\gamma$. Intuitively, to explore the local curvature information of $F(\cdot)$, we can devise a tighter local surrogate $G_{k,e_{k}/t_{k}}(\cdot)$ with $\bm{\mu}\preceq_{K}e_{k}\preceq_{K}\bm{\ell}$. In this case, the linear convergence rate of Algorithm 6 can be further improved by using a tighter local surrogate $G_{k,e_{k}/t_{k}}(\cdot)$.
To narrow the surrogate gap and better capture the local curvature information, we compute $e_{k}$ by Barzilai-Borwein method, namely, we set
$e_{k}:=\frac{\left\langle{JF(x^{k})-JF(x^{k-1}),x^{k}-x^{k-1}}\right\rangle}{\left\lVert{x^{k}-x^{k-1}}\right\rVert^{2}}.$ (15)
Lemma 5.3.
Assume that $F(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$, let $\{x^{k}\}$ be the sequence generated by Algorithm 6, where $e_{k}$ is defined as in (15). Then, the following statements hold:
  • (i)
    $\bm{\mu}\preceq_{K}e_{k}\preceq_{K}\bm{\ell}$;
  • (ii)
    $t_{k}\geq\min_{c^{*}\in C_{e}}\{\gamma\left\langle{c^{*},e_{k}}\right\rangle/\left\langle{c^{*},\bm{\ell}}\right\rangle\}$;
Proof.
Assertion (i) follows by the strong $k$-convexity and $K$-smoothness of $F(\cdot)$, and the definition of $e_{k}$. We can obtain the assertion (ii) by using the similar arguments as in the proof of Proposition 5.1. ∎
6 Descent methods for VOPs with polyhedral cones
In this section, we consider the VOPs that $K$ is a polyhedral cone with nonempty interior. Without loss of generality, for the polyhedral cone $K$, there exists a transform matrix $A\in\mathbb{R}^{l\times m}$ with $m\leq l$ such that
$K:=\{x\in\mathbb{R}^{m}:0\preceq Ax\}.$
In this case, for any $a,b\in\mathbb{R}^{m}$, $a\preceq_{K}b$ can be equivalently represented as $Aa\preceq Ab$. Denote $A_{i}$ the $i$-th row vector of $A$. For the polyhedral cone $K$, we denote the set of transform matrices as follows:
$\mathcal{A}:=\left\{A\in\mathbb{R}^{l\times m}:AK=\mathbb{R}^{l}_{+}\right\}.$
Notably, in practice, the polyhedral cone $K$ is often defined by a specific transform matrix $A$. This raises a crucial question: does the transform matrix $A$ affect the performance of descent methods for VOPs with polyhedral cones?
6.1 Steepest descent method for VOPs with polyhedral cones
By using the transform matrix $A\in\mathcal{A}$, the steepest descent direction subproblem for VOPs with polyhedral cones is formulated as follows:
$\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{\lambda\in\Delta_{l}}\left\langle{\lambda,AJF(x^{k})d}\right\rangle+\frac{1}{2}\left\lVert{d}\right\rVert^{2}.$ (16)
In the next subsection, we claim that the preceding subproblem can be efficiently solved via its duality. The complete $K$-steepest descent method for VOPs with polyhedral cones is described as follows:
Data: $x^{0}\in\mathbb{R}^{n},\gamma\in(0,1)$
1 Select a transform matrix $A\in\mathcal{A}$
2 for $k=0,1,...$ do
3       Update $d^{k}$ as the minimizer of (16)
4       if $d^{k}=0$ then
5             return $C$-stationary point $x^{k}$
6      else
7            Compute the stepsize $t_{k}\in(0,1]$ in the following way:
$t_{k}:=\max\left\{\gamma^{j}:j\in\mathbb{N},A(F(x^{k}+\gamma^{j}d^{k})-F(x^{k}))\preceq\gamma^{j}\left(AJF(x^{k})d^{k}+\frac{1}{2}\left\lVert{d^{k}}\right\rVert^{2}e\right)\right\},$
where $e:=(1,...,1)^{T}$
8             $x^{k+1}:=x^{k}+t_{k}d^{k}$
9       end if
10
11 end for
Algorithm 7 K-steepest_descent_method_for_VOPs_with_polyhedral_cones
Lemma 6.1.
Assume that $F(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$, where $K=\{x\in\mathbb{R}^{m}:0\preceq Ax\}.$ Let $\{x^{k}\}$ be the sequence generated by Algorithm 7. Then, the following statements hold:
  • (i)
    $t_{k}\geq\min_{i\in[l]}\{\gamma/\left\langle{A_{i},\bm{\ell}}\right\rangle\}$;
  • (ii)
    $\{x^{k}\}$ converges to an efficient solution $x^{*}$ of (VOP);
  • (iii)
    $\left\lVert{x^{k+1}-x^{*}}\right\rVert\leq\sqrt{1-\min_{i\in[l]}t_{k}\left\langle{A_{i},\bm{\mu}}\right\rangle}\left\lVert{x^{k}-x^{*}}\right\rVert,~{}\forall k\geq 0$.
Proof.
The assertions can be obtained by using the similar arguments as in the proof of Proposition 5.1 and Lemma 4.1. ∎
Remark 6.1.
The subproblem (16) can be reformulated as
$\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{c^{*}\in C}\left\langle{c^{*},JF(x^{k})d}\right\rangle+\frac{1}{2}\left\lVert{d}\right\rVert^{2},$
where $C:=conv\{A_{i},i\in[l]\}$ is a base of $K^{*}$. In other words, selecting a transform matrix in (16) is equivalent to selecting a base of $K^{*}$. Consequently, the linear convergence rate of Algorithm 7 is sensitive to the choice of $A$. When $K=\mathbb{R}^{m}_{+}$, the steepest descent method [17] fixs $A=I_{m}$, i.e., $C=\Delta_{m}$.
6.2 Barzilai-Borwein descent method for VOPs with polyhedral cones
In general, for the $e_{k}$ defined as (15) we have $e_{k}\preceq_{K}\bm{\ell}$, which can be written as $Ae_{k}\preceq A\bm{\ell}$. We denote $\alpha^{k}\in\mathbb{R}^{l}_{++}$ as follows:
$\alpha_{i}^{k}=\left\{\begin{aligned} &\max\left\{\alpha_{\min},\min\left\{\frac{\langle s_{k-1},y^{k-1}_{i}\rangle}{\left\lVert{s_{k-1}}\right\rVert^{2}},\alpha_{\max}\right\}\right\},&\langle s_{k-1},y^{k-1}_{i}\rangle&>0,\\ &\max\left\{\alpha_{\min},\min\left\{\frac{\left\lVert{y^{k-1}_{i}}\right\rVert}{\left\lVert{s_{k-1}}\right\rVert},\alpha_{\max}\right\}\right\},&\langle s_{k-1},y^{k-1}_{i}\rangle&<0,\\ &\alpha_{\min},&\langle s_{k-1},y^{k-1}_{i}\rangle&=0,\end{aligned}\right.$ (17)
for all $i\in[l]$, where $s_{k-1}=x^{k}-x^{k-1},\ y^{k-1}_{i}$ is the $i$-th row vector of $A(JF(x^{k})-JF(x^{k-1}))$, $\alpha_{\max}$ is a sufficient large positive constant and $\alpha_{\min}$ is a sufficient small positive constant. The Barzilai-Borwein descent direction is defined as the minimizer of
$\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{\lambda\in\Delta_{l}}\left\langle{\lambda,AJF(x^{k})d+\frac{1}{2}\left\lVert{d}\right\rVert^{2}\alpha^{k}}\right\rangle.$ (18)
Alternatively, we use the similar strategy in Algorithm 4, the Barzilai-Borwein descent direction with appropriate base is defined as the minimizer of
$\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{\lambda\in\Delta^{\alpha^{k}}_{l}}\left\langle{\lambda,AJF(x^{k})d}\right\rangle+\frac{1}{2}\left\lVert{d}\right\rVert^{2},$ (19)
where $\Delta^{\alpha^{k}}_{l}:=\{c^{*}\in\mathbb{R}^{l}_{+}:\left\langle{c^{*},\alpha^{k}}\right\rangle=1\}$. The subproblem can be reformulated as follows:
$\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{\lambda\in\Delta_{l}}\left\langle{\lambda,\Lambda^{k}AJF(x^{k})d}\right\rangle+\frac{1}{2}\left\lVert{d}\right\rVert^{2},$ (20)
where
$\Lambda^{k}:=\begin{bmatrix}\frac{1}{\alpha^{k}_{1}}&&\\ &\ddots&\\ &&\frac{1}{\alpha^{k}_{l}}\end{bmatrix}.$
By optimality conditions, the minimizer of (20) can be written as
$d^{k}=-(\Lambda^{k}AJF(x^{k}))^{T}\lambda^{k},$
where $\lambda^{k}\in\Delta_{l}$ is a solution of the following dual problem:
$\min\limits_{\lambda\in\Delta_{l}}\frac{1}{2}\left\|(\Lambda_{k}AJF(x^{k}))^{T}\lambda\right\|^{2}.$ (DP)
Remark 6.2.
The dual problem (DP) is lower dimensional quadratic programming with unit simplex constraint (the vertices of unit simplex constraint are known), then it can be solved by Frank-Wolfe/conditional gradient method efficiently [see, e.g., 41, 7]. However, dual problem of (18) reads as
$\min\limits_{\lambda\in\Delta_{l}}\frac{1}{2}\frac{\left\|(AJF(x^{k}))^{T}\lambda\right\|^{2}}{\sum\limits_{i\in[l]}\lambda_{i}\alpha^{k}_{i}},$
which is not easy to solve.
Remark 6.3.
If $K=\mathbb{R}^{m}_{+}$ and $A=I_{m}$, the subproblem (20) reduces to that of the BBDMO [7]. Consequently, transforming (18) into (20) gives a new insight into BBDMO. Comparing the subproblems (16) and (19), we conclude that the differences between the directions of the steepest descent and the Barzilai-Borwein descent lie in the choice of base of $K^{*}$.
The complete $K$-Barzilai-Borwein descent method for VOPs with polyhedral cones is described as follows:
Data: $x^{0}\in\mathbb{R}^{n},\gamma\in(0,1)$
1 Choose $x^{-1}$ in a small neighborhood of $x^{0}$
2 for $k=0,1,...$ do
3       Update $\alpha^{k}$ as (17)
4       Update $d^{k}$ as the minimizer of (20)
5       if $d^{k}=0$ then
6             return $C$-stationary point $x^{k}$
7      else
8            Compute the stepsize $t_{k}\in(0,1]$ in the following way:
$t_{k}:=\max\left\{\gamma^{j}:j\in\mathbb{N},A(F(x^{k}+\gamma^{j}d^{k})-F(x^{k}))\preceq\gamma^{j}\left(AJF(x^{k})d^{k}+\frac{1}{2}\left\lVert{d^{k}}\right\rVert^{2}\alpha^{k}\right)\right\}$
9             $x^{k+1}:=x^{k}+t_{k}d^{k}$
10       end if
11
12 end for
Algorithm 8 K-Barzilai-Borwein_descent_method_for_VOPs_with_polyhedral_cones
Lemma 6.2.
Assume that $F(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$, where $K=\{x\in\mathbb{R}^{m}:0\preceq Ax\}.$ Let $\{x^{k}\}$ be the sequence generated by Algorithm 8. Then, the following statements hold:
  • (i)
    $A\bm{\mu}\preceq\alpha^{k}\preceq A\bm{\ell}$;
  • (ii)
    $t_{k}\geq\min_{i\in[l]}\{\gamma\alpha^{k}_{i}/\left\langle{A_{i},\bm{\ell}}\right\rangle\}$;
  • (iii)
    $\{x^{k}\}$ converges to an efficient solution $x^{*}$ of (VOP);
  • (iv)
    $\left\lVert{x^{k+1}-x^{*}}\right\rVert\leq\sqrt{1-\min_{i\in[l]}{t_{k}\left\langle{A_{i},\bm{\mu}}\right\rangle}/{\alpha^{k}_{i}}}\left\lVert{x^{k}-x^{*}}\right\rVert,~{}\forall k\geq 0$.
Proof.
The assertions can be obtained by using the similar arguments as in the proof of Proposition 5.1 and Lemma 4.1. ∎
A large stepsize may speed up the convergence of Algorithm 8. Accordingly, the Armijo line search can be applied, namely, compute the stepsize $t_{k}\in(0,1]$ in the following way:
$t_{k}:=\max\left\{\gamma^{j}:j\in\mathbb{N},A(F(x^{k}+\gamma^{j}d^{k})-F(x^{k}))\preceq\sigma\gamma^{j}AJF(x^{k})d^{k}\right\},$ (21)
where $\sigma\in(0,1)$.
The following results show that the convergence rate of Algorithm 8 is not sensitive to the choice of transform matrix. More specifically, the descent direction $d^{k}$ and stepsize $t_{k}$ of Algorithm 8 are invariant for some $A\in\mathcal{A}$.
Proposition 6.1.
(Affine Invariance) Let $A^{1},A^{2}\in\mathcal{A}$, $d^{k}_{1},t_{k}^{1}$ and $d^{k}_{2},t_{k}^{2}$ be the descent directions and stepsize generated by Algorithm 8 with $A^{1}$ and $A^{2}$, respectively. If $\alpha_{\min}<\alpha_{i}^{k,1},\alpha_{i}^{k,2}<\alpha_{\max}$, $i\in[l]$, we have $d^{k}_{1}=d^{k}_{2}$ and $t_{k}^{1}=t_{k}^{2}$.
Proof.
Denote $A_{i}^{1}$ and $A_{i}^{2}$ the the $i$-th row vector of $A^{1}$ and $A^{2}$, respectively. Before presenting the main results, we rewritten the subproblem (20) as follows:
$\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{i\in[l]}\left\langle{\frac{A_{i}}{\alpha^{k}_{i}},JF(x^{k})d}\right\rangle+\frac{1}{2}\left\lVert{d}\right\rVert^{2}.$
Recall that $A^{1},A^{2}\in\mathcal{A}$, there exists a vector $a\in\mathbb{R}^{l}_{++}$ such that
$\left\{A^{1}_{i}:i\in[l]\right\}=\left\{a_{i}A^{2}_{i}:i\in[l]\right\}.$ (22)
We claim the following assertion:
$\left\{\frac{A^{1}_{i}}{\alpha_{i}^{k,1}}:i\in[l]\right\}=\left\{\frac{A^{2}_{i}}{\alpha_{i}^{k,2}}:i\in[l]\right\}.$ (23)
This, together with the reformulated subproblem, implies that $d^{k}_{1}=d^{k}_{2}$. Therefore, $t_{k}^{1}=t_{k}^{2}$ is a consequence of (23). Next, we prove that assertion (23) holds. For any $i\in[l]$, it follows by (22) that there exist $j\in[l]$ such that $A^{1}_{i}=a_{j}A^{2}_{j}$. Notice that $\alpha_{\min}<\alpha_{i}^{k,1},\alpha_{i}^{k,2}<\alpha_{\max}$, $i\in[l]$, without loss of generality, we assume that $\alpha_{i}^{k,1}={\left\lVert{A^{1}_{i}(JF(x^{k})-JF(x^{k-1}))}\right\rVert}/{\left\lVert{s_{k-1}}\right\rVert}$ and $\alpha_{j}^{k,2}={\left\lVert{A^{2}_{j}(JF(x^{k})-JF(x^{k-1}))}\right\rVert}/{\left\lVert{s_{k-1}}\right\rVert}$. Then,
$\frac{A^{1}_{i}}{\alpha_{i}^{k,1}}=\frac{A^{1}_{i}\left\lVert{s_{k-1}}\right\rVert}{\left\lVert{A^{1}_{i}(JF(x^{k})-JF(x^{k-1}))}\right\rVert}=\frac{A^{2}_{j}\left\lVert{s_{k-1}}\right\rVert}{\left\lVert{A^{2}_{j}(JF(x^{k})-JF(x^{k-1}))}\right\rVert}=\frac{A^{2}_{j}}{\alpha_{j}^{k,2}},$
where the second equality is due to the fact that $A^{1}_{i}=a_{j}A^{2}_{j}$. Thus, we have
$\left\{\frac{A^{1}_{i}}{\alpha_{i}^{k,1}}:i\in[l]\right\}\subseteq\left\{\frac{A^{2}_{i}}{\alpha_{i}^{k,2}}:i\in[l]\right\}.$
The relation
$\left\{\frac{A^{2}_{i}}{\alpha_{i}^{k,2}}:i\in[l]\right\}\subseteq\left\{\frac{A^{1}_{i}}{\alpha_{i}^{k,1}}:i\in[l]\right\}$
follows the similar arguments, this concludes the proof. ∎
Remark 6.4.
Assume $\hat{\mathcal{A}}\subset\mathcal{A}$ is bounded and $A_{i}$ is bounded away from 00 for all $A\in\hat{\mathcal{A}}$. Then, there exists $\alpha_{\min}$ and $\alpha_{\max}$ such that the assumption $\alpha_{\min}<\alpha_{i}^{k}<\alpha_{\max}$, $i\in[l]$ holds for all $A\in\hat{\mathcal{A}}$ with $\left\langle{s_{k-1},{A_{i}(JF(x^{k})-JF(x^{k-1}))}}\right\rangle\neq 0$. For the case with linear objective, $\left\langle{s_{k-1},{A_{i}(JF(x^{k})-JF(x^{k-1}))}}\right\rangle=0$ may hold. As illustrated in [7, Example 3], for any $A_{1},A_{2}\in\hat{\mathcal{A}}$, we have $d^{k}_{1}\approx d^{k}_{2}$ with sufficient small $\alpha_{\min}$. As a result, we conclude that the performance of Algorithm 8 is not sensitive to the choice of $A$.
7 Second-order methods from a majorization-minimization perspective
Let $K=\mathbb{R}^{m}_{+}$, Fliege et al. [16] proposed a Newton method for MOPs, the Newton direction is defined as the optimal solution of
$\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{\lambda\in\Delta_{m}}\left\langle{\lambda,JF(x^{k})d+\frac{1}{2}d^{T}D^{2}F(x^{k})d}\right\rangle,$ (24)
where $d^{T}D^{2}F(x^{k})d:=(d^{T}\nabla^{2}F_{1}(x^{k})d,...,d^{T}\nabla^{2}F_{m}(x^{k})d)$. The objective-wise quadratic approximation serves as good surrogate in Newton method but makes the subproblem challenging to solve. From a majorization-minimization perspective, the surrogate should strike a balance between better approximation and lower computational cost. In the remainder of this section, we develop an efficient second-order method. Interestingly, the proposed method coincides with the preconditioned Barzilai-Borwein method for MOPs [9].
Firstly, we note that the challenging nature of the subproblem (24) arises primarily from selecting $\Delta_{m}$ as the base for +$\mathbb{R}^{m}_{+}$. By instead choosing a general base in the subproblem:
$\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{\lambda\in\Delta^{\alpha^{k}}_{m}}\left\langle{\lambda,JF(x^{k})d+\frac{1}{2}d^{T}D^{2}F(x^{k})d}\right\rangle,$ (25)
where $\Delta^{\alpha^{k}}_{m}:=\{c^{*}\in\mathbb{R}^{m}_{+}:\left\langle{c^{*},\alpha^{k}}\right\rangle=1\}$, it can be reformulated as follows:
$\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{\lambda\in\Delta_{m}}\sum\limits_{i\in[m]}\lambda_{i}\frac{\left\langle{\nabla F_{i}(x^{k}),d}\right\rangle}{\alpha^{k}_{i}}+\frac{1}{2}\sum\limits_{i\in[m]}\lambda_{i}\frac{\left\langle{d,\nabla^{2}F_{i}(x^{k})d}\right\rangle}{\alpha^{k}_{i}}.$ (26)
This raises a direct question: how to select an appropriate $\alpha^{k}$? Intuitively, if
$\frac{\left\langle{d^{k},\nabla^{2}F_{i}(x^{k})d^{k}}\right\rangle}{\alpha^{k}_{i}}=\sum\limits_{i\in[m]}\lambda^{k}_{i}\frac{\left\langle{d^{k},\nabla^{2}F_{i}(x^{k})d^{k}}\right\rangle}{\alpha^{k}_{i}}$
for $i\in[m]$, by setting $B_{k}\approx\sum_{i\in[m]}\lambda^{k}_{i}{\nabla^{2}F_{i}(x^{k})}/{\alpha^{k}_{i}}$, we reformulate (26) into the following subproblem:
$\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{\lambda\in\Delta_{m}}\sum\limits_{i\in[m]}\lambda_{i}\frac{\left\langle{\nabla F_{i}(x^{k}),d}\right\rangle}{\alpha^{k}_{i}}+\frac{1}{2}\left\lVert{d}\right\rVert^{2}_{B_{k}},$ (27)
where $\alpha^{k}_{i}\approx\left\langle{d^{k},\nabla^{2}F_{i}(x^{k})d^{k}}\right\rangle/\left\lVert{d^{k}}\right\rVert^{2}_{B_{k}}$ for $i\in[m]$. Since $d^{k}$ is unknown before solving the subproblem, we estimate it with $d^{k-1}$ and set
$\alpha^{k}_{i}=\frac{\left\langle{\nabla F_{i}(x^{k})-\nabla F_{i}(x^{k-1}),x^{k}-x^{k-1}}\right\rangle}{\left\lVert{x^{k}-x^{k-1}}\right\rVert^{2}_{B_{k}}}$
for $i\in[m]$. Unfortunately, it is important to note that $\lambda^{k}$ is unavailable prior to computing the subproblem, and $\alpha^{k}$ is determined based on $B_{k}$. As an alternative, we can approximate $B_{k}$ using $\sum_{i\in[m]}\lambda^{k-1}_{i}{\nabla^{2}F_{i}(x^{k})}/{\alpha^{k-1}_{i}}$ instead of $\sum_{i\in[m]}\lambda^{k}_{i}{\nabla^{2}F_{i}(x^{k})}/{\alpha^{k}_{i}}$. Consequently, by updating $B_{k}$ via the BFGS method, the subproblem (27) coincides with that of preconditioned Barzilai-Borwein method for MOPs, which can be efficiently solved using the Frank-Wolfe method [9].
Remark 7.1.
Chen et al. [9] provided a detailed explanation of why the combination of the Barzilai-Borwein (BB) method and the BFGS method is essential in the multi-objective optimization context, particularly from a preconditioning perspective. Specifically, these two approaches function as complementary preconditioners for the search direction. While their effects tend to overlap in single-objective optimization, in the multi-objective setting, they fulfill distinct roles and demonstrate improved performance when applied together. In this section, we also highlight the indispensability of the Barzilai-Borwein rule in the quasi-Newton method from a majorization-minimization perspective.
8 Numerical Results
In this section, we present numerical results to demonstrate the performance of Barzilai-Borwein descent methods for VOPs (BBDVO) with polyhedral cones. We also compare BBDVO with steepest descent method for VOPs (SDVO) and equiangular direction method [28] for VOPs (EDVO). All numerical experiments were implemented in Python 3.7 and executed on a personal computer with an Intel Core i7-11390H, 3.40 GHz processor, and 16 GB of RAM. For all tested algorithms, we used Armijo line search (21) with $\sigma=10^{-4}$ and $\gamma=0.5$. To ensure that the algorithms terminate after a finite number of iterations, for all tested algorithms we used the stopping criterion $\|d(x)\|\leq 10^{-6}$. We also set the maximum number of iterations to 500. The test algorithms were executed on several test problems, and the problem illustration is given in Table 1. The dimensions of variables and objective functions are presented in the second and third columns, respectively. $x_{L}$ and $x_{U}$ represent lower bounds and upper bounds of variables, respectively. For each problem, we used the same initial points for different tested algorithms. The initial points were randomly selected within the specified lower and upper bounds. Dual subproblems of different algorithms were efficiently solved by Frank-Wolfe method. The recorded averages from the 200 runs include the number of iterations, the number of function evaluations, and the CPU time.
Table 1: Description of all test problems used in numerical experiments
Problem $n$ $m$ $x_{L}$ $x_{U}$ Reference
BK1 2 2 (-5,-5) (10,10) [24]
DD1 5 2 (-20,…,-20) (20,…,20) [11]
Deb 2 2 (0.1,0.1) (1,1) [12]
FF1 2 2 (-1,-1) (1,1) [24]
Hil1 2 2 (0,0) (1,1) [23]
Imbalance1 2 2 (-2,-2) (2,2) [7]
JOS1a 50 2 (-2,…,-2) (2,…,2) [26]
LE1 2 2 (-5,-5) (10,10) [24]
PNR 2 2 (-2,-2) (2,2) [39]
WIT1 2 2 (-2,-2) (2,2) [46]
For the tested problems, the partial order are induced by polyhedral cones +$\mathbb{R}^{2}_{+}$, $K_{1}$, and $K_{2}$, respectively, where
$K_{1}:=\{x\in\mathbb{R}^{2}:5x_{1}-x_{2}\geq 0,~{}-x_{1}+5x_{2}\geq 0\}\subseteq\mathbb{R}^{2}_{+},$
and
$K_{2}:=\{x\in\mathbb{R}^{2}:5x_{1}+x_{2}\geq 0,~{}x_{1}+5x_{2}\geq 0\}\supseteq\mathbb{R}^{2}_{+}.$
8.1 Numerical results for VOPs with $K=\mathbb{R}^{2}_{+}$
In this case, we denote the set of transform matrices $\mathcal{A}_{0}:=\{A:A\mathbb{R}^{2}_{+}=\mathbb{R}^{2}_{+}\}$. For SDVO, we choose $A^{0}=I_{2}\in\mathcal{A}_{0}$ and $\hat{A}^{0}\in\mathcal{A}_{0}$ in subproblem, respectively, where $\hat{A}^{0}:=\{A:A_{i}=A^{0}_{i}/\max\{1,\left\lVert{\nabla F_{i}(x^{0})}\right\rVert_{\infty}\},~{}i=1,2\}$ [^1] The scale strategy is initially proposed in [21] due to numerical reasons.. For EDVO, normalization is applied for each of gradients in the transformed subproblem, which implies that EDVO is also not sensitive to the choice of transform matrix. As a result, we choose $A=A^{0}$ in subproblems of EDVO and BBDVO.
Table 2: Number of average iterations (iter), number of average function evaluations (feval), and average CPU time (time($ms$)) of SDVO, EDVO and BBDVO implemented on different test problems with $K=\mathbb{R}^{2}_{+}$
Problem SDVO with $A=A^{0}$ SDVO with $A=\hat{A}^{0}$ EDVO with $A=A^{0}$ BBDVO with $A=A^{0}$
iter feval time iter feval time iter feval time iter feval time
BK1 1.00 2.00 0.21 1.00 2.00 0.39 27.87 32.11 2.37 1.00 1.00 0.23
DD1 71.09 199.63 12.14 43.14 163.92 10.77 354.26 354.26 33.48 7.49 8.91 1.26
Deb 57.86 376.40 13.95 55.14 350.18 16.50 10.18 26.02 1.42 4.41 6.58 0.84
FF1 28.95 29.11 3.97 16.00 61.75 4.40 9.07 10.00 0.93 4.91 6.13 0.78
Hil1 25.04 82.89 5.24 12.87 49.58 4.34 17.68 45.45 2.74 11.32 12.15 2.08
Imbalance1 88.23 178.31 11.86 387.91 388.32 40.39 405.13 405.92 34.90 2.61 3.54 0.46
JOS1a 311.23 311.23 34.89 16.53 18.05 3.47 65.67 65.67 7.87 1.00 1.00 0.29
LE1 22.44 52.57 3.40 7.98 29.78 1.70 11.34 29.26 1.35 4.55 7.03 0.81
PNR 11.33 48.80 2.13 12.23 15.97 1.76 13.65 21.89 1.41 4.18 4.74 0.66
WIT1 73.45 462.47 16.43 139.82 146.50 14.22 10.06 16.60 1.07 3.53 3.62 0.55
8.2 Numerical results for VOPs with $K=K_{1}$
[…]
Table 3: Number of average iterations (iter), number of average function evaluations (feval), and average CPU time (time($ms$)) of SDVO, EDVO and BBDVO implemented on different test problems with $K=K_{1}$
Problem SDVO with $A=A^{1}$ SDVO with $A=\hat{A}^{1}$ EDVO with $A=A^{1}$ BBDVO with $A=A^{1}$
iter feval time iter feval time iter feval time iter feval time
BK1 1.00 2.65 0.23 7.49 9.85 1.18 43.66 47.15 4.04 1.00 1.00 0.16
DD1 93.30 376.98 18.65 118.15 120.10 14.63 495.62 495.62 48.61 42.07 47.34 6.61
Deb 102.47 878.36 29.21 27.88 146.52 6.47 68.61 145.74 8.80 35.69 71.88 6.55
FF1 33.20 105.30 6.25 17.24 56.85 3.45 40.00 57.24 4.35 16.09 17.03 2.51
Hil1 29.77 159.99 7.65 20.13 90.95 5.06 21.78 52.25 3.37 17.66 18.27 2.97
Imbalance1 77.34 309.32 13.53 397.37 397.44 44.40 492.52 492.52 42.95 28.78 31.18 4.49
JOS1a 81.13 81.13 9.69 6.57 17.06 2.53 106.12 106.12 12.63 1.00 1.00 0.24
LE1 14.02 69.28 3.01 12.37 55.82 2.45 15.63 39.36 1.89 6.29 7.47 1.09
PNR 30.34 178.44 6.61 13.10 38.79 2.38 19.95 25.23 2.05 9.57 10.77 1.65
WIT1 245.72 1095.05 71.75 198.45 245.46 22.39 405.71 408.04 38.08 151.35 156.72 21.66
8.3 Numerical results for VOPs with $K=K_{2}$
[…]
Table 4: Number of average iterations (iter), number of average function evaluations (feval), and average CPU time (time($ms$)) of SDVO, EDVO and BBDVO implemented on different test problems with $K=K_{2}$
Problem SDVO with $A=A^{2}$ SDVO with $A=\hat{A}^{2}$ EDVO with $A=A^{2}$ BBDVO with $A=A^{2}$
iter feval time iter feval time iter feval time iter feval time
BK1 25.16 96.73 3.31 6.67 11.79 1.14 19.66 25.82 1.97 1.00 1.00 0.24
DD1 18.90 69.25 3.83 44.21 45.34 4.90 120.74 128.80 12.21 4.86 5.16 0.79
Deb 19.48 164.15 6.58 18.30 135.99 4.89 22.44 143.05 4.89 9.44 48.91 2.21
FF1 13.54 19.00 2.19 14.75 67.65 3.22 17.57 125.12 4.49 4.78 5.59 0.94
Hil1 15.40 79.01 4.08 17.99 84.51 4.54 15.74 39.73 2.47 8.24 9.09 1.51
Imbalance1 26.43 111.38 4.62 500.00 610.90 55.78 500.00 500.04 44.23 4.36 5.77 0.71
JOS1a 53.38 53.38 5.99 4.65 11.66 2.18 42.96 42.96 5.36 1.00 1.00 0.31
LE1 9.64 47.62 2.22 14.81 79.35 3.06 10.12 49.15 1.81 7.55 42.97 1.73
PNR 18.77 114.16 4.18 13.49 51.13 2.38 10.75 34.71 1.49 6.63 8.64 1.02
WIT1 24.11 147.56 5.89 308.99 973.42 44.31 10.07 28.11 1.42 8.42 9.98 1.34
$K=\mathbb{R}^{2}_{+}$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$K=K_{1}$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
$K=K_{2}$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) BK1 (b) DD1 (c) FF1 (d) Imbalance1 (e) WIT1
Figure 1: Numerical results in value space obtained by BBDVO for problems BK1, DD1, FF1, Imbalance1 and WIT1 with $K=\mathbb{R}^{2}_{+}$, $K=K_{1}$ and $K=K_{2}$, respectively.
For test problems with different partial orders, the number of average iterations (iter), number of average function evaluations (feval), and average CPU time (time(ms)) of the different algorithms are listed in Tables 2, 3 and 4, respectively. We conclude that BBDVO outperforms SDVO and EDVO, especially for problems DD1 and Imbalance1. For SDVO, its performance is sensitive to the choice of transform matrix, changing the transform matrix in subproblem cannot improve the performance on all test problems. Naturally, a question arises that how to choose an appropriate transform matrix for a specific test problem in SDVO. It is worth noting that BBDVO can be viewed as SDVO with variable transform matrices ($\Lambda^{k}A$ is a transform matrix of $K$) and thus enjoys promising performance on these test problems. This provides a positive answer to the question. EDVO can also be viewed as SDVO with variable transform matrices, it generates descent directions with norm less than $1$ (the minimizer of subproblem is the minimal norm element of the convex hull of some unit vectors), decelerating the convergence in large-scale problems (the initial point may be far from the Pareto set).
Fig. 1 plots the final points obtained by BBDVO on problems BK1, DD1, FF1, Imbalance1 and WIT1 with $K=\mathbb{R}^{2}_{+}$, $K=K_{1}$ and $K=K_{2}$, respectively. We can observe that enlarging the partial order cone reduces the number of obtained Pareto critical points, especially in the long tail regions, where improving one objective function slightly can sacrifice the others greatly. As a result, we can use an order cone containing the non-negative orthant in real-world MOPs to obtain Pareto points with a better trade-off.
9 Conclusions
In this paper, we develop a unified framework and convergence analysis of descent methods for VOPs from a majorization-minimization perspective. We emphasize that the convergence rate of a descent method can be improved by narrowing the surrogate functions. By changing the base in subproblems, we elucidate that choosing a tighter surrogate function is equivalent to selecting an appropriate base of the dual cone. From the majorization-minimization perspective, we employ Barzilai-Borwein method to narrow the local surrogate functions and propose a Barzilai-Borwein descent method for VOPs polyhedral cone. The proposed method is not sensitive to the choice of transform matrix, which affects the performance of SDVO. Numerical experiments confirm the efficiency of the proposed method.
From a majoriztion-minimization perspective, we also rediscover the preconditioned Barzilai-Borwein method for MOPs. This highlights the versatility of the majorization-minimization principle as a powerful framework for designing novel algorithms in vector optimization. In future work, it is worth analyzing proximal gradient methods and high-order methods for VOPs within the majorization-minimization framework. Additionally, exploring solution methods for VOPs with non-polyhedral cones presents an intriguing avenue for further research.
Acknowledgement
This work was funded by the Major Program of the National Natural Science Foundation of China [grant numbers 11991020, 11991024]; the Key Program of the National Natural Science Foundation of China [grant number 12431010]; the General Program of the National Natural Science Foundation of China [grant number 12171060]; NSFC-RGC (Hong Kong) Joint Research Program [grant number 12261160365].
Data Availability
The data that support the findings of this study are available from the first author, [Jian Chen], upon reasonable request.